### v1

# RAISE: Enhancing Scientific Reasoning in LLMs via Step-by-Step Retrieval

Abstract
Scientific reasoning requires not only long-chain reasoning processes, but also knowledge of domain-specific terminologies and adaptation to updated findings. To deal with these challenges for scientific reasoning, we introduce RAISE, a step-by-step retrieval-augmented framework which retrieves logically relevant documents from in-the-wild corpus. RAISE is divided into three steps: problem decomposition, logical query generation, and logical retrieval. We observe that RAISE consistently outperforms other baselines on scientific reasoning benchmarks. We analyze that unlike other baselines, RAISE retrieves documents that are not only similar in terms of the domain knowledge, but also documents logically more relevant.
RAISE: Enhancing Scientific Reasoning in LLMs via
Step-by-Step Retrieval
Minhae Oh1, Jeonghye Kim2, Nakyung Lee1, Donggeon Seo3, Taeuk Kim3, Jungwoo Lee1, 1Seoul National University, 2KAIST, 3Hanyang University {minhaeoh, junglee,lnktop}@snu.ac.kr, jeonghye.kim@kaist.ac.kr, {donggunseo,kimtaeuk}@hanyang.ac.kr
1 Introduction
Recently, large language models (LLMs) have demonstrated remarkable capabilities for diverse scientific domains that require intricate reasoning and the integration of diverse knowledge sources (Zhang et al., 2024c, a; Prabhakar et al., 2025; Rueda et al., 2025). In contrast to many other domains, scientific reasoning is especially demanding due to the need for advanced reasoning skills, knowledge of specialized terminology, and awareness of continually evolving knowledge. To enhance the capability of LLMs to solve scientific reasoning problems, two approaches are widely used: (1) step-wise reasoning, and (2) retrieval-augmented generation. For step-wise reasoning, it has been shown that solving reasoning problems through step-wise reasoning process is more effective than attempting to solve them in a single step, similar to how humans approach challenging problems Wei et al. (2022); Zhao et al. (2024a); Jin et al. (2024); Lai et al. (2024); Yao et al. (2023). For retrieval-augmented generation (RAG), it is known that retrieving relevant external knowledge can mitigate hallucination and deal with timely knowledge, which is especially important for science domain Lewis et al. (2020b); Bang et al. (2025); Zhong et al. (2025); Xiong et al. (2024).
To utilize the benefits of both worlds, recent works have adopted RAG for step-wise reasoning on various domains (Jeong et al., 2024; Zhang et al., 2025a). However, most of the works focus on relatively simple multi-hop question answering tasks or non-scientific domains (Guan et al., 2025; Jin et al., 2025; Wang et al., 2025) or assume task-related retrieval corpus which is hard to obtain compared to in-the-wild retrieval source such as Wikipedia (Zhong et al., 2025; Qian et al., 2023; Anand et al., 2023). Solving challenging scientific reasoning tasks such as graduate-level biology, physics, etc, using an in-the-wild corpus is very challenging since merely retrieving superficial knowledge is insufficient for these tasks. Instead, the retrieved information should contain relevant logics needed to solve the problem (Rueda et al., 2025). Moreover, the knowledge required for each intermediate step can vary significantly even within the same problem. Without considering the evolving information needed for each reasoning process, RAG might even deteriorate the downstream task performance. The question of what to search for and how to retrieve the appropriate external knowledge for each step when solving scientific reasoning tasks is underexplored.
Figure 1: Overview of RAISE. RAISE is divided into three steps: (1) Problem Decomposition, (2) Logical Query Generation, and (3) Logical Retrieval. Given an scientific reasoning question, we decompose the problem into multiple subquestions and search queries. Then, we generate a logical query that contains logic-enriched contexts. Finally, we retrieve relevant documents from in-the-wild corpus for each subquestion and condition the documents to assist the step-by-step reasoning for scientific reasoning problems.
To address these challenges, we introduce RAISE (step-by-step Retrieval-Augmented Inference for Scientific rEasoning), a retrieval-augmented framework tailored for step-wise scientific reasoning. Our framework is divided into three steps: (1) problem decomposition, (2) logical query generation, and (3) logical retrieval. For problem decomposition, LLMs decompose the original question into subquestions along with search queries that can assist in solving each reasoning step. Then, for each search query, LLMs convert the search query into a logical query, which is a query that contains logical explanations. Since logical queries contain more information about what external knowledge is needed to solve the problem, they enable retrievers to retrieve based on logic similarity rather than superficial domain similarity. Finally, we perform retrieval separately for each step, obtaining documents that are logically relevant to the particular step. Instead of assuming task-relevant or well-curated retrieval source, such as question-answer pool of relevant domains, we retrieve from in-the-wild source such as Wikipedia, which enables applying to challenging real-world scenarios.
We evaluate our framework on three challenging scientific reasoning benchmarks: GPQA Rein et al. (2024), SuperGPQA Team et al. (2025), and MMLU Hendrycks et al. (2021); Wang et al. (2024b).RAISE consistently outperforms various baselines approaches that include applying RAG or problem decomposition. This indicates that our framework effectively retrieves relevant documents that are needed to solve scientific reasoning tasks step-by-step. Through further analysis, we observe that RAISE retrieves more logically relevant documents for each step compared to baselines, highlighting that logical similarity is crucial for these challenging reasoning tasks in addition to domain similarity. Our code is available at https://github.com/minhaeoh/RAISE.
2 Preliminary
Step-by-Step Reasoning in LLMs.    LLMs are capable of performing multi-step reasoning over complex input queries by internally chaining intermediate inferences. This step-by-step reasoning process involves decomposing a question into sub-problems, maintaining coherence across steps, and generating a final answer. Formally, given a query $x$, the model implicitly constructs a latent reasoning trajectory $\{r_{t}\}_{t=1}^{T}$, and generates the answer $y$ conditioned on this chain:
$p(y\mid x)=\sum_{r_{1},\dots,r_{T}}p(y\mid r_{1:T},x)\cdot\prod_{t=1}^{T}p(r_{t}\mid r_{<t},x).$
However, standard LLMs rely solely on their parametric knowledge, which limits performance in scenarios requiring up-to-date or external information.
RAG for Single-Step Reasoning.    We address the task of generating a response $y$ given an input $x$, enhanced by retrieval from an external corpus $\mathcal{D}$. RAG combines a retriever and a generator to condition the output on both the input and relevant documents.
A standard language model defines:
$p(y\mid x)=\prod_{t=1}^{T}p(y_{t}\mid y_{<t},x).$
In RAG, generation is conditioned on retrieved documents $\{d_{j}\}_{j=1}^{k}$, typically approximated as:
$p(y\mid x)\approx\sum_{j=1}^{k}p(y\mid x,d_{j})\cdot p(d_{j}\mid x).$
The retriever encodes queries and documents via $f_{q}(x)$ and $f_{d}(d)$, scoring relevance by:
$\text{sim}(x,d)=f_{q}(x)^{\top}f_{d}(d).$
Top-$k$ documents are retrieved, and a generator (e.g., BART Lewis et al. (2020a), T5 Raffel et al. (2020)) produces $y$ based on both $x$ and $d_{i}$.
Figure 2: Examples comparing query generation methods (Step-back, HyDE, and RAISE) for the same subquestion.
3 RAISE
We propose RAISE (step-by-step Retrieval-Augmented Inference for Scientific rEasoning), a retrieval-augmented generation framework for scientific reasoning designed to support multi-step reasoning through fine-grained, step-aware retrieval. Rather than retrieving documents based on the entire problem as a single query, RAISE decomposes the problem into intermediate reasoning steps and performs targeted retrieval for each. The overview of RAISE is provided in Figure 1 and Algorithm 1. The method consists of three main stages:
  1. (1)
    Problem Decomposition, where the original question $x$ is broken down into subquestions $r_{1},\dots,r_{n}$ along with corresponding search queries $q_{1},\dots,q_{n}$ that are intended to retrieve relevant knowledge for each step;
  2. (2)
    Logical Query Generation, where each search query $q_{i}$ is further enriched into a logic aware explanation $q_{i}^{*}$, which serves as the final retrieval query; and
  3. (3)
    Logical Retrieval, where external knowledge $D_{i}$ is retrieved for each subquestion $r_{i}$ and used to generate the subanswer $a_{i}$.
Problem Decomposition.    RAISE begins with the decomposition of a scientific reasoning problem $x$ into a sequence of intermediate subquestions $\{r_{i}\}_{i=1}^{n}$, each corresponding to a distinct reasoning step. This planning process is performed by prompting a language model $P_{\theta}$ with $x$ and prompt $p_{1}$ to identify the logical structure required to solve the problem. Formally, we generate:
$\{(r_{i},q_{i})\}_{i=1}^{n}\sim P_{\theta}(\cdot\mid x,p_{1})$ (1)
While recent work has increasingly explored problem decomposition, many retrieval-based approaches still operate at the problem level, retrieving information using a single query without explicitly modeling intermediate reasoning steps. In contrast, this stage produces a structured sequence of subquestions that reflects the multi-step nature of scientific reasoning.
Along with each subquestion $r_{i}$, a corresponding search query $q_{i}$ is generated to guide retrieval. These queries are not used directly for retrieval but rather serve as an initial query for the next stage. As a result, this stage outputs subquestion-query pairs $\{(r_{i},q_{i})\}_{i=1}^{n}$, forming the basis for step-wise retrieval and generation.
Logical Query Generation.    In the second stage, each initial search query $q_{i}$ and its corresponding subquestion $r_{i}$ are jointly used to generate a logically enriched logical query $q_{i}^{*}$. Search queries $q_{i}$ are short key phrases, where the logical process is absent. Therefore, using $q_{i}$ for retrieval will focus on domain similarity instead of logical similarity. Instead, this step constructs a logically coherent query that aligns with the reasoning intent of the subquestion. The model is prompted with both $q_{i}$ and $r_{i}$, along with a reformulation prompt $p_{2}$ (e.g. add essential scientific or mathematical explanation the search result would contain):
$q_{i}^{*}\sim P_{\theta}(\cdot\mid q_{i},r_{i},p_{2})$ (2)
Using only the subquestion $r_{i}$ to generate a query often results in noisy or overly specific formulations that include problem-specific variables or conditions irrelevant to general knowledge retrieval. On the other hand, using only the initial query $q_{i}$ lacks sufficient context to capture the intended reasoning path. By combining both, we enable the model to synthesize a logical query that is purpose-driven and aligned with the knowledge required to answer $r_{i}$. Even if the reformulated query $q_{i}^{*}$ contains facutal inaccuracies, it tends to retrieve passages from corpus $\mathcal{C}$ that are logically relevant and supportive of the reasoning required for solving the original problem. Figure 2 presents example queries generated by RAISE ,Step-back and HyDE, illustrating RAISE’s ability to generate logical queries that are well-aglined with the reasoning intent.
Logical Retrieval.    The reformulated query $q_{i}^{*}$ is used to retrieve external knowledge:
$D_{i}=\mathcal{R}(q_{i}^{*},\mathcal{C},k)$ (3)
Here, $\mathcal{R}$ is a top-$k$ retrieval function over in-the-wild corpus $\mathcal{C}$ (e.g., Wikipedia), and $D_{i}=\{d_{i1},\dots,d_{ik}\}$ denotes the retrieved documents. The goal is not to retrieve direct answers but to obtain passages that contain relevant background knowledge or reasoning procedures necessary to solve $r_{i}$. Also, since irrelevant documents might be also retrieved, we set a retrieval threshold $T$, where we retrieve documents that have higher retrieval similarity than $T$.
After retrieval, for each subquestion $r_{i}$, the model predicts its solution $a_{i}$. Importantly, the prediction is conditioned not only on the retrieved documents $D_{i}$, but also on the original question $x$ and the previously answered subquestions $\{(r_{j},a_{j})\}_{j=1}^{i-1}$. We write:
$a_{i}\sim P_{\theta}(\cdot\mid x,\{(r_{j},a_{j})\}_{j=1}^{i-1},r_{i},D_{i},p_{3})$ (4)
Finally, all subanswers are aggregated to predict the final answer $y$:
$y\sim P_{\theta}(\cdot\mid x,\{(r_{i},a_{i})\}_{i=1}^{n},p_{4})$ (5)
This step-wise approach ensures that the model constructs its final answer based on grounded and explicitly structured intermediate reasoning.
Table 1: Comparison of various reasoning strategies across GPQA, SuperGPQA, and MMLU. RAISE consistently outperforms other approaches for scientific reasoning benchmarks.
Direct Reasoning Decomposed Reasoning
CoT PS Step-back CoT+RAG Least-to-Most Step-back RAG RAISE
GPQA
Overall 42.42 39.39 43.94 45.96 44.95 44.44 45.95 51.01
SuperGPQA
science-hard 7.54 9.05 9.05 7.54 6.03 5.03 6.03 10.05
science-middle 15.08 14.07 17.59 12.56 14.57 15.08 15.08 19.60
engineering-hard 7.54 4.02 5.03 7.54 10.05 7.54 8.04 10.55
MMLU
(Pro) Chemistry 25.44 26.50 26.86 25.18 24.56 22.70 22.97 28.36
(STEM) College Chemistry 49.50 45.00 44.00 43.00 45.40 43.00 46.00 51.00
4 Experiments
4.1 Experimental Setup
Datasets.    To evaluate our method in settings that require multi-step scientific reasoning, we select three datasets focused on high-level STEM and professional reasoning. We provide more details on the evaluation dataset in Appendix A.
  • GPQA (Graduate-level Physics Question Answering) Rein et al. (2024): A collection of graduate-level physics, biology, and chemistry questions requiring multi-step reasoning and access to external scientific fact. We use this dataset as our primary benchmark.
  • SuperGPQA (Team et al., 2025): This dataset evaluates the knowledge and reasoning abilities of Large Language Models across 285 graduate-level disciplines. We focus on three subsets that are challenging enough and require step-by-step reasoning in the science domain: science-hard, science-middle, and engineering-hard.
  • MMLU (Hendrycks et al., 2021; Wang et al., 2024b): Among various domains and difficulty levels of MMLU (Hendrycks et al., 2021) and MMLU-Pro (Wang et al., 2024b), we focus on chemistry domain since it requires step-by-step reasoning with knowledge of specialized terminology.
Retriever Configuration.    We adopt Dense Passage Retrieval (DPR) Karpukhin et al. (2020) as our retriever, using the publicly available model trained on the Natural Questions (NQ) dataset Kwiatkowski et al. (2019).
Language Model.    For GPQA, which is the primary focus of our study, we use Mistral Small 3.1-Instruct-2503 Mistral AI (2025), a 24B state-of-the-art small-scale open-sourced LLM that achieves strong performance on scientific reasoning tasks. For the remaining tasks, we use the smaller LLaMA 3.1-8B model Grattafiori et al. (2024) due to computational constraints, applying it to SuperGPQA, and MMLU. We provide detailed hyperparameters and experiment setups in Appendix B.
4.2 Baselines
To assess the importance of multi-step reasoning and step-aware retrieval, we conduct experiments with two groups of baselines: Direct Reasoning and Decomposed Reasoning. Direct Reasoning methods solve the problem in a single inference step without decomposing it into subquestions. This group includes three prompt-based approaches and one retrieval-augmented method. The three prompting methods-Chain-of-Thought (CoT) Wei et al. (2022), Plan-and-Solve (PS) Wang et al. (2023), Step-back Zheng et al. (2024) -differ in how the prompt guides the language models’ reasoning. Additionally, we include CoT+RAG Lewis et al. (2020b), a conventional retrieval-augmented method that retrieves documents using the original question as query.
Decomposed Reasoning methods first break down the original problem into subquestions and solve them sequentially. This group includes Least-to-Most Zhou et al. (2023), which solves each subquestion directly without retreival, Step-back, which applies principle extraction at the subquestion level, and RAG, which retrieves and uses evidence for each subquestion individually. More detailed descriptions of each baseline can be found in Appendix B.1
4.3 Main Results
As shown in Table 1, our proposed method, RAISE, consistently outperforms all baseline reasoning strategies across three benchmark datasets: GPQA, SuperGPQA, and MMLU. On GPQA, RAISE achieves 51.01%, surpassing the next-best method (CoT+RAG) by over 5 percentage points. On SuperGPQA, which consists of especially challenging science and engineering questions, RAISE achieves the highest performance across all subdomains, with gains especially notable in the science-middle subset (19.60%, +2 pts over the best baseline). On MMLU, RAISE also shows significant improvement, particularly in (Pro) Chemistry (28.36%) and STEM College Chemistry (51.00%), indicating its robustness across both professional and academic scientific reasoning tasks. This result shows that RAISE effectively decomposes scientific reasoning questions and retrieves documents that help the model to solve the questions.
Unlike RAISE, decomposed reasoning methods do not always lead to better performance. This is often due to a lack of logically relevant knowledge especially for smaller LLMs, which limits their ability to reason effectively across multiple steps (Hosseini et al., 2024; Xu et al., 2025). In fact, the original papers of our baseline models that use decomposed reasoning mostly focused on proprietary large LLMs while we focus on open-sourced smaller LLMs. Despite well-structured decomposition into subquestions, a lack of logically relevant knowledge may cause the model to produce hallucinations or unsupported reasoning steps. Especially for smaller models that lack sufficient prior knowledge, problem decomposition can become counterproductive unless logically relevant documents are retrieved to support the reasoning process.
Also, we observe that in some cases, models without retrieval outperform their retrieval-augmented counterparts. While retrieval is generally expected to enhance performance by providing additional information, these results reveal that naively applying retrieval can be ineffective or even harmful. This issue becomes particularly pronounced in settings involving open-domain or in-the-wild corpora, where the risk of retrieving noisy or irrelevant content is high. Conventional RAG methods often retrieve documents that are only superficially related or worse, focus on irrelevant variables or distractors rather than the core principles required for reasoning. These findings highlight the importance of retrieving logically relevant knowledge, and further underscore the effectiveness of our method.
Table 2: Comparison of decomposed reasoning strategies with retrieval. RAISE consistently outperforms baselines across datasets.
GPQA SuperGPQA MMLU
RAG 45.95 9.72 24.67
Step-back+RAG 43.43 10.05 33.03
HyDE 28.28 9.22 35.99
RAISE 51.01 13.57 39.68
4.4 Additional Results
Impact of Logical Query Generation.    To evaluate the impact of logical query generation in retrieval-augmented reasoning, we conduct an experiment comparing our method, RAISE, with three RAG-based decomposed reasoning baselines: RAG, step-back+RAG (Zheng et al., 2024), and HyDE (Gao et al., 2023a). While all methods incorporate problem decomposition and retrieval, they differ in how the retrieval query is constructed.
RAISE begins with a short key phrase known as the initial query, and expands it into a logical query that captures not only domain similarity but also the underlying reasoning intent of the subquestion. In contrast, the standard RAG baseline directly uses the subquestion itself as the retrieval query. The step-back+RAG approach attempts to abstract a general principle from the subquestion and uses that principle as the query. HyDE generates a synthetic passage in response to the subquestion and uses this generated text as the retrieval query, aiming to enrich the query context.
As shown in Table 2, RAISE outperforms all other RAG-based baselines across all benchmarks. These results confirm that RAISE’s logical query generation goes beyond retrieving documents that are merely domain-relevant or superficially similar, enabling the retrieval of knowledge that is logically aligned with the reasoning required to solve the problem. We conduct further analysis in Section 4.5. For SuperGPQA and MMLU, Table 2 reports average scores across subsets; full per-subset results are provided in the Appendix 5.
Figure 3: Performance comparison between RAISE-Direct and RAISE across datasets.
Table 3: Evaluation on GPQA with various LLMs with different scales: LLaMA 3.1-8B, GPT-4o mini, and Mistral Small 3.1. RAISE consistently shows the best performance across all settings.
LLaMA GPT Mistral
CoT 22.22 40.91 42.42
PS 27.78 38.90 39.39
Step-back 28.28 40.90 43.94
CoT+RAG 23.23 40.40 45.96
Least-to-Most 26.26 45.45 44.95
Step-back (Decomp.) 28.28 42.42 44.44
RAG 24.24 42.93 45.95
RAISE 30.30 47.98 51.01
Impact of Problem Decomposition.    To further assess the importance of problem decomposition, we also evaluate a variant of our method that omits this step and directly performs logical query generation and retrieval without breaking the problem into subquestions, as shown in Figure 3. This version, referred to as RAISE-Direct, showed lower performance compared to the full version of RAISE. These results indicate that problem decomposition plays a critical role in guiding the retrieval process and structuring the reasoning pathway.
Applying RAISE to various LLMs.    To assess the generalizability of RAISE across different LLM scales, we evaluate its performance on GPQA using LLaMA 3.1-8B (Grattafiori et al., 2024) and GPT-4o mini (OpenAI, 2024), in addition to Mistral(used in our main experiments). As shown in Table 3, RAISE demonstrates consistent improvements over other baselines, exhibiting a similar trend to our main results with Mistral-24B. This shows that the effect of RAISE is not limited to a specific type of LLM, but can be applied to various LLMs with different scales.
Table 4: Examples where RAISE-retrieved documents provide logically relevant information for scientific reasoning compared to baseline RAG retrieval.
Subquestion RAG RAISE Explanation
What is the product of the reaction of 2,8-dimethylspiro[4.5] decan-6-ol with sulfuric acid? Carbylamine reaction … synthesis of an isocyanide by the reaction of a primary amine, chloroform, and base. The alkene acts as a nucleophile and attacks the proton, following Markovnikov’s rule. In the second step, an HO molecule bonds to the more substituted carbon… The RAISE-retrieved document explains the acid-catalyzed dehydration mechanism of alcohols, directly aligning with the transformation of 2,8-dimethylspiro[4.5]decan-6-ol to a ketone.
What is the concentration of OH ions in a solution of 0.3 M Ba(OH)2? Normality is an ambiguous measure of the concentration of a solution. It needs a definition of the equivalence factor… Barium hydroxide is a chemical compound with the formula Ba(OH)2(H2O). Barium hydroxide can be prepared by dissolving BaO in water… The Ba centers adopt a square anti-prismatic geometry. The RAISE-retrieved document clearly identifies barium hydroxide as $\mathrm{Ba(OH)_{2}}$ and explains its dissociation behavior in water, directly supporting the calculation of $[\mathrm{OH}^{-}]$ concentration.
How does the scalar potential (V) change when considering the retarded time $t_{r}$? Compensating variation in economics … additional money needed after a price change. The calculation of the retarded time $t_{r}=t-|\vec{r}-\vec{r^{\prime}}|/c$ is a speed-distance-time relation. The field observed at time $t$ must have been emitted at this earlier time. The RAISE-retrieved document directly explains the concept of retarded time in electrodynamics as the time delay due to the finite propagation speed of electromagnetic fields, which is essential for deriving the scalar potential $V(\vec{r},t)$ of a moving charge.
4.5 Analysis of RAISE
To further investigate our hypothesis that RAISE retrieves documents that are logically more relevant compared to other baselines, we use LLM-as-a-judge (GPT-4o-mini) to evaluate the logical relevancy of the retrieved documents. Conditioned on the question, subquestion for a specific step, and the retrieved documents, the evaluator model evaluates the logical relevancy among 4 levels of logical relevancy: (1) Not relevant at all, (2) Superficially relevant (topically related but logically unhelpful), (3) Partially logically relevant (some useful reasoning content), and (4) Fully logically relevant (logically sufficient to solve the subquestion).
The results are illustrated in Figure 4. Compared to other baselines that also applies RAG, RAISE has the lowest ratio of documents that are irrelevant at all or only superficially relevant (relevant in terms of domain knowledge, but not relevant logically) and highest ratio of documents that are at least partially logically relevant. This indicates that RAISE avoids retrieving documents that may interrupt the reasoning process for scientific reasoning through logical query generation.
We also qualitatively demonstrate the examples when RAISE retrieves logically relevant documents compared to convential RAG in Table 4. While RAG often retrieves documents that are topically related yet fail to address the reasoning needs of the subquestion, RAISE consistently identifies documents that include essential scientific principles, mechanisms, or equations. For instance, in questions involving chemical reactions or physical formulas, RAISE surfaces materials that explain the specific reactivity or the retarded time calculation, whereas RAG may return vague definitions or unrelated economic concepts. These cases illustrate how RAISE’s retrieval is not only domain-aware but also aligned with the logical demands of solving complex scientific problems.
Figure 4: Logical Relevancy of Retrieved Documents. Unlike other baselines, RAISE has higher ratio of documents that are logically relevant and lower ratio of documents that are irrelevant or superficially relevant.
5 Related Works
LLMs for Scientific Reasoning.    Recent works have shown that LLMs can be applied for challenging scientific reasoning tasks. Unlike other domains, scientific reasoning requires not only step-by-step thinking, but also knowledge of specialized terminology and adaptation to continually evolving knowledge. Due to this challenging nature, many benchmarks have been proposed recently to tackle scientific reasoning with LLMs (Rein et al., 2024; Zhong et al., 2025; Zhang et al., 2025b; Wang et al., 2024a; Jiang et al., 2024). Many works enhance scientific reasoning capabilities of LLMs through domain-specific training (Taylor et al., 2022; Prabhakar et al., 2025; Zhang et al., 2024b), step-by-step reasoning (Rueda et al., 2025; Rein et al., 2024), or retrieval of external knowledge or tools (Ma et al., 2024; Zhong et al., 2025; Wellawatte et al., 2025; Li et al., 2025). Unlike previous works, we focus on applying step-by-step document retrieval from in-the-wild corpus without assuming access to well-curated and domain-specific corpus.
Step-wise Reasoning.    A growing body of research has shown that decomposing complex problems into structured intermediate steps can enhance the reasoning abilities of LLMs. An influential early approach, Chain-of-Thought prompting Wei et al. (2022), introduced explicit, sequential reasoning steps, making the model’s thought process more transparent and coherent. This inspired methods such as Plan-and-Solve Wang et al. (2023), which emphasizes high-level planning before answering, and Step-Back Prompting Zheng et al. (2024), which encourages abstraction by prompting the model to reflect before solving. Least-to-Most prompting Zhou et al. (2023) extends this by breaking down tasks into simpler subproblems, solved in increasing order of difficulty.
While prior work has focused on prompting strategies that help LLMs better use their internal reasoning capabilities, our work addresses a complementary challenge: enabling LLMs to retrieve and apply information from in-the-wild sources like Wikipedia, particularly during step-wise problem solving. We investigate how external evidence can be integrated at each step to improve reasoning beyond what internal knowledge alone can achieve.
Retrieval-Augmented Generation.    Retrieval-Augmented Generation (RAG) Lewis et al. (2020b) was initially proposed to improve LLMs’ factual accuracy and knowledge by retrieving relevant external documents during generation Lewis et al. (2020b); Karpukhin et al. (2020); Izacard and Grave (2021); Khattab and Zaharia (2020).
Recently, RAG has been extended for multi-hop reasoning, performing retrieval iteratively at multiple reasoning steps Liu et al. (2025); Zhao et al. (2024b); Zheng et al. (2024). In parallel, query reformulation and expansion techniques have been developed to enhance retrieval. Instead of using the original question, models generate enriched queries through prompting, such as intermediate answers or summaries. For example, HyDE Gao et al. (2023a) and CSQE Lei et al. (2024) demonstrate that carefully crafted queries greatly improve retrieval in complex, multi-step tasks.
Building on this line of work, we redesign query expansion techniques with the specific goal of retrieving documents that contain the key logic or underlying principles required at each step of a step-wise reasoning process. This enables the model to supplement its limited internal knowledge with external sources, leading to more complete problem solving, especially in complex, multi-step tasks.
6 Conclusion
We introduce RAISE, a step-by-step retrieval framework for scientific reasoning. We first decompose the problem into multiple subquestions and search queries, and then generate logical queries and retrieve logically relevant documents from in-the-wild corpus. We demonstrate the effectiveness of RAISE on three scientific reasoning benchmarks by comparing with various baselines. Our analysis shows that RAISE retrieves documents that are not only relevant in terms of the domain (e.g. definition of specificalized terminology) but also logically relevant documents for each subquestion, assisting the step-by-step reasoning process required for scientific reasoning.
Limitations
In our study, we primarily conducted experiments using relatively smaller language models, mainly due to computational resource constraints. As a result, the performance observed may not fully reflect the potential of larger-scale models. Future work could explore scaling up the model size to assess its impact on reasoning capabilities and overall performance.
For the retrieval process, we used only DPR as the retriever. While DPR has been widely recognized for its effectiveness, we believe that more recent retrieval models Gao et al. (2023b); Basu et al. (2024); Shi et al. (2023), could lead to a substantial improvement in performance. Incorporating these advanced retrievers remains a promising direction for future research.
Furthermore, our work is particularly focused on scientific reasoning tasks. While this domain presents unique challenges and complexities, extending our approach to other domains such as legal, commonsense, or medical reasoning could be an interesting and valuable avenue for further investigation. Such cross-domain studies may help validate the generalizability and robustness of RAISE.
Additionally, RAISE relies on multiple calls to large language models during the logical query generation phase. Although this multi-step querying is essential for generating effective logical queries that maximize retrieval efficiency, exploring strategies to reduce the number of LLM invocations, such as more efficient prompting, intermediate caching, or lightweight planning modules, could enhance the overall efficiency and practicality of our method in real-world applications.
Appendix A Dataset Details
GPQA Rein et al. (2024)    This dataset consists of physics, biology, and chemistry questions written by domain experts. We use GPQA diamond subset, which consist of 198 high-quality questions selected based on human performance. Specifically, this subset includes questions that both experts answer correctly while the majority of non-experts fail to solve. Each question typically demands multi-step reasoning, precise formula manipulation, and access to external scientific facts (e.g., physical constants, definitions). Due to its alignment with our target setting, GPQA serves as the primary evaluation benchmark throughout our experiments.
For GPQA, the original dataset does not include standardized multiple-choice labeled as (A), (B), (C) and (D). To ensure consistency during evaluation, we proprocessed each question by randomly shuffling the correct answer along with the three distractors, and assigning them uniformly to choice labels (A) through (D).
SuperGPQA (Team et al., 2025)    SuperGPQA is a large-scale benchmark designed to evaluate graduate-level reasoning across 13 disciplines, 72 fields, and 285 graduate-level disciplines. In alignment with the scientific reasoning focus of our work, we select science and engineering domains for evaluation. Each domain is further divided by three difficulty levels(easy, medium, and hard). To reduce computational overhead while maintaining consistency, we randomly sample 199 questions per subset using a fixed seed (42). Specifically, our experiments include 199 examples each from science-hard, science-middle, and engineering-hard subsets.
MMLU (Hendrycks et al., 2021; Wang et al., 2024b)    The MMLU benchmark covers a wide range of subjects across multiple domains. For out experiments, we focus on STEM and Professional categories. The STEM contains university-level science and engineering subjects such as college mathematics and computer science, while the Professional category covers specialized fields that typically require professional training or advanced education, including law, medicine, and chemistry. We specifically select two subsets: college chemistry from MMLU-STEM and professional chemistry from MMLU-Pro. These subsets are chosen to evaluate our method’s ability to perform scientific reasoning in both academic and professional contexts involving complex domain knowledge.
Appendix B Experiment Details
B.1 Baseline Details
CoT Wei et al. (2022); Kojima et al. (2023) We apply Chain-of-Thought prompting for direct reasoning, where the model is encouraged to explicitly generate intermediate reasoning steps through prompting (Think step by step).
PS Wang et al. (2023) Plan-and-Solve prompting is a direct reasoning strategy where the model first outlines a plan and then executes it step by step. We use the same prompt as the original paper.
Step-back (Direct reasoning) Zheng et al. (2024) Step-back prompting guides reasoning using abstracted principles derived from the original question. In our implementation, we follow the original Step-back paper’s prompting strategy to extract high-level principles necessary for solving the question. These extracted principles are then provided together with the original question as input to guide the model’s reasoning process.
CoT+RAG Lewis et al. (2020b) We implement CoT+RAG by combining Chain-of-Thought prompting with retrieval, where the model is prompted to solve the problem step-by-step while also leveraging external knowledge. Specifically, we provide the model with a CoT-style prompt encouraging step-by-step reasoning, alongside the original question and documents retrieved using the original question as the search query.
Least-to-Most Zhou et al. (2023) Least-to-Most is a decomposed reasoning strategy that breaks down a complex problem into a sequence of simpler subquestions, which are then solved sequentially without retrieval augmentation. This subquestion decomposition pipeline serves as the foundational structure for other decomposed reasoning methods as well.
Step-back (Decomposed reasoning) Zheng et al. (2024) We implement Step-back for decomposed reasoning by applying the Step-back prompting method to each subquestion in a decomposed reasoning framework. While the original Step-back paper does not cover the application of this method to decomposed subquestions, we extend it for a fair comparison with our approach. Specifically, after decomposing the original question into subquestions, we use the Step-back prompting strategy to extract a high-level principle for each subquestion, and then provide the subquestion along with its corresponding principle to guide the model’s reasoning.
RAG (Decomposed reasoning) Liu et al. (2025) We implement RAG by first decomposing the original problem into subquestions and then retrieving documents using each subquestion as a query. The retrieved documents are provided to the model along with the corresponding subquestion to support its reasoning.
B.2 Retriever Configuration
We use the pre-trained DPR encoder from the ’facebook/dpr-question_encoder-single-nq-base’ model Karpukhin et al. (2020), which is a BERT-based encoder trained for open-domain question answering. This encoder is trained on the Natural Question (NQ) dataset Kwiatkowski et al. (2019) and is designed to map questions into 768-dimensional dense vector representations for retrieval.
For the retrieval corpus, we use the preprocessed Wikipedia passages provided by ’facebook/wiki_dpr’ Karpukhin et al. (2020), a corpus widely used to evaluate DPR-based retrieval models. This corpus is constructed from the December 20, 2018 Wikipedia dump, where each article is split into multiple, disjoint text blocks of 100 words, resulting in approximately 21 million passages. Each passage is accompanied by the title of the wikipedia page it comes from along with DPR embedding.
To enable efficient retrieval over the passage embeddings, we use an exact FAISS index. FAISS (Facebook AI Similarity Search) Johnson et al. (2019); Douze et al. (2024) is widely used library for fast similarity search over dense vectors.
Throughout all experiment, we retrieve top-10 documents per query.
B.3 RAISE
To reduce the impact of potentially irrelevant documents by DPR, we apply a similarity threshold $T$ in RAISE.Specifically, we discard any retrieved passage whose DPR similarity score falls below $T$. DPR similarity is computed as the inner product between L2-normalized query and passage embeddings. Higher scores indicate greater semantic similarity, with values closer to 1 representing stronger alignment between the query and passage.
We set $T=0.84$ for GPQA, SuperGPQA, and MMLU-Pro, which are composed of more challenging reasoning problems. For MMLU-STEM (college chemistry), we use a slightly lower threshold of $T=0.80$, considering that the questions are generally simpler than those in other datasets.
B.4 Algorithm
Input: Original question $x$, prompts $\mathcal{P}=\{p_{1},p_{2},p_{3},p_{4}\}$, corpus $\mathcal{C}$
Output: Final answer $y$
Step 1: Problem Decomposition
Generate subquestions and initial queries:
$\{(r_{i},q_{i})\}_{i=1}^{n}\sim P_{\theta}(\cdot\mid x,p_{1})$
for $i=1$ to $n$ do
       Step 2: Logical Query Generation
       Reformulate initial query:
        $q_{i}^{*}\sim P_{\theta}(\cdot\mid r_{i},q_{i},p_{2})$
      Step 3: Knowledge Retrieval
       Retrieve top-$k$ documents:
        $D_{i}=\mathcal{R}(q_{i}^{*},\mathcal{C},k)$
      Step 4: Subquestion Answering
       if $i=1$ then
              $a_{i}\sim P_{\theta}(\cdot\mid x,r_{1},D_{1},p_{3})$
      else
              $a_{i}\sim P_{\theta}(\cdot\mid x,\{(r_{j},a_{j})\}_{j=1}^{i-1},r_{i},D_{i},p_{3})$
       end if
end for
Step 5: Final Answer Composition
Generate final answer using all subanswers:
$y\sim P_{\theta}(\cdot\mid x,\{(r_{i},a_{i})\}_{i=1}^{n},p_{4})$
Algorithm 1 RAISE Inference Procedure
B.5 Prompts
B.5.1 Baseline Prompts
Prompt for CoT You are solving a multiple choice question. Think step by step and show your reasoning clearly. At the end, state your answer in the format: "The final answer is (X)". Here, X must be the correct letter choice. Question: [Problem here] Answer Choices: [Answer choices here] Solution:
Prompt for Plan-and-Solve Let’s first understand the problem and devise a plan to solve the problem. Then, let’s carry out the plan and solve the problem step by step. Use the following format exactly: <Plan> Step 1: [state what you will do] Step 2: [state next reasoning step] … (as needed) <Solution> Step 1: [carry out Step 1] Step 2: [carry out Step 2] … (as needed) "The final answer is (X)." Here, X must be the correct letter choice. Question: [Problem here] Answer Choices: [Answer choices here] Solution:
Prompt for Step-back principle abstraction You are an expert at Science. You are given a Science problem. Your task is to extract the Science concepts and principles involved in solving the problem. What are the principles behind this question? End your response with "End of generation" after you answer the instructions. Question: [Subquestion here] Principles Involved:
Prompt for Step-back You are an expert at Science. You are given a Science problem and a set of principles involved in solving the problem. Solve the problem step by step by folowing the principles. At the end, state your answer in the format: "The final answer is (X)". Here, X must be the correct letter choice. Question: [Problem here] Principles: [Principles here] Answer Choices: [Answer choices here] Solution:
Prompt for hyDE query generation Generate a paragraph that answers the question. End your response with "End of generation" after you answer the instructions. Question: [Subquestion here] Explanation:
B.5.2 RAISE Prompts
Prompt for Problem Decomposition You are given a multiple-choice question. Break this problem into essential subquestions that directly help solve the original problem. Each subquestion MUST also include its search query. Each search query should reflect scientific or mathematical knowledge needed to answer the subquestion. STRICT FORMAT REQUIREMENTS: 1. For each subquestion, you MUST provide exactly two parts in this order: - The subquestion - A search query for that subquestion 2. Use EXACTLY this format for each subquestion: Subquestion 1: [your specific subquestion] Search Query for Subquestion 1: [Write a search query someone might realistically use to learn how to answer this subquestion] Question: [Problem here] Answer Choices: [Answer choices here]
Prompt for Logical Query Generation You are given a subquestion and a search query. The search query is a realistic phrase that someone might use to find knowledge or reasoning support to answer the subquestion. Your task is to anticipate what essential scientific or mathematical explanation the search result would contain, and write it concisely (2–3 sentences). Focus only on the core concept or principle that would help answer the subquestion. Avoid restating the subquestion, and do not include unrelated or overly general information. Subquestion: [Subquestion resulting from Problem Decomposition] Search Query: [Search query resulting from Problem Decomposition] Explanation:
Prompt for Solving Subquestions with Documents You are solving a multiple-choice question. Question is decomposed into several subquestions. You will be given: 1. The original multiple-choice question 2. Previous subquestions and their solutions (if any) 3. The current subquestion to solve 4. Documents that are relevant to the current subquestion Your task: - Carefully read the original question, any previous subquestion and their solutions, and the current subquestion - Use the information from the retrieved documents to solve the current subquestion - Also use your existing knowledge to solve the current subquestion - Your solution should be detailed and logically structured Documents: [Retrieved document] Question: [Problem here] Answer Choices: [Answer choices here] Previous subquestions and their solutions: [Previously generated subquestions and solutions] Current subquestion to solve: Subquestion [Step num]: [Subquestion] Subquestion [Step num] Solution:
Prompt for Generating Final Answer You are solving a multiple-choice question. Question is decomposed into several subquestions. Each subquestion has already been solved. Your task is to carefully read the original question and the several subquestion solutions, then use them to determine the final answer. Think step by step and then finish your answer with "The final answer is (X)" where X is the correct letter choice. Original Question: Question: [Problem here] Answer Choices: [Answer choices here] Subquestions and Solutions: [Generated stepwise subproblems and solutions] Final Solution:
Prompt for Evaluation with GPT You are given the following three items: - Original Problem: [Problem here] - Subquestion: [Subquestion here] - Retrieved Document: [Document here] Your task is to evaluate how helpful the retrieved document is for answering the subquestion. Please follow these instructions: - Do not just check if the topic is related. - Instead, check if the document includes information that helps someone reason through and solve the subquestion. - Focus on whether the document supports actual thinking or steps needed to get the answer. Give your final judgment using only one of the following ratings: - "No relevance at all" - does not have any domain similarity - "Superficially relevant" - has domain similarity (only superficially) but does not have any logical relevance to the subquestion. For example, the document might mention the same topic as the subquestion, but it does not provide any information that helps solve the subquestion. - "Partially relevant" - has domain similarity and has some logical relevance to the subquestion. For example, the document might provide some information that helps solve the subquestion, but it does not provide all the logical steps needed. - "Fully relevant" - has domain similarity and has almost all logical relevance to the subquestion. For example, the document provides enough relevant logical steps to solve the subquestion. Then explain your reasoning briefly. Output Format: Helpfulness Rating: <one of the 4 options above> Explanation: <your short explanation>
Appendix C Additional Results
Figure 5: Domain-wise accuracy comparison between CoT and RAISE on the GPQA Diamond subset
Figure 5 shows the domain-wise accuracy on the GPQA Diamond dataset. We compare the performance of RAISE against Chain-of-Thought (CoT) prompting across all domains. RAISE outperforms or matches CoT in nearly all domains, with only one domain where CoT shows higher accuracy. These results demonstrate RAISE’s robustness and its ability to generalize across diverse areas of graduate-level scientific reasoning.
Table 5 provides the full subset results for each benchmark, complementing the averaged scores shown in Table 2. Across all datasets, RAISE consistently outperforms other retrieval-augmented baselines. The consistent improvement across subsets supports our central hypothesis that constructing queries enriched with logical intent is essential for retrieving logically relevant knoweldge.
Table 5: Comparison of decomposed reasoning strategies with retrieval. Subset-level results for each benchmark, complementing the averaged scores reported in the main text.
RAG Step-back +RAG HyDE RAISE
GPQA
Overall 45.95 43.43 48.48 51.01
SuperGPQA
science-hard 6.03 5.53 7.54 10.55
science-middle 15.08 15.58 13.07 19.60
engineering-hard 8.04 9.05 7.04 10.55
MMLU
(Pro) Chemistry 22.53 23.06 22.97 28.36
(STEM) College Physics 46.00 43.00 49.00 51.00

### v2

@@
 Scientific reasoning requires not only long-chain reasoning processes, but also knowledge of domain-specific terminologies and adaptation to updated findings.
-To deal with these challenges for scientific reasoning, we introduce RAISE, a step-by-step retrieval-augmented framework which retrieves logically relevant documents from in-the-wild corpus.
+To address these challenges in scientific reasoning, we introduce RAISE, a step-by-step retrieval-augmented framework which retrieves logically relevant documents from an in-the-wild corpus.
 RAISE is divided into three steps: problem decomposition, logical query generation, and logical retrieval.
@@
 Step-by-Step Retrieval
-Minhae Oh1, Jeonghye Kim2, Nakyung Lee1, Donggeon Seo3, Taeuk Kim3, Jungwoo Lee1, 1Seoul National University, 2KAIST, 3Hanyang University {minhaeoh, junglee,lnktop}@snu.ac.kr, jeonghye.kim@kaist.ac.kr, {donggunseo,kimtaeuk}@hanyang.ac.kr
+Minhae Oh1, Jeonghye Kim2, Nakyung Lee1, Donggeon Seo3, Taeuk Kim3, Jungwoo Lee1, 1Seoul National University, 2KAIST, 3Hanyang University
 1 Introduction
-Recently, large language models (LLMs) have demonstrated remarkable capabilities for diverse scientific domains that require intricate reasoning and the integration of diverse knowledge sources (Zhang et al., 2024c, a; Prabhakar et al., 2025; Rueda et al., 2025).
+Recently, large language models (LLMs) have demonstrated remarkable capabilities in scientific domains that require intricate reasoning and the integration of diverse knowledge sources Zhang et al.
+(2024c, a); Prabhakar et al.
+(2025); Rueda et al.
+(2025).
 In contrast to many other domains, scientific reasoning is especially demanding due to the need for advanced reasoning skills, knowledge of specialized terminology, and awareness of continually evolving knowledge.
 To enhance the capability of LLMs to solve scientific reasoning problems, two approaches are widely used: (1) step-wise reasoning, and (2) retrieval-augmented generation.
-For step-wise reasoning, it has been shown that solving reasoning problems through step-wise reasoning process is more effective than attempting to solve them in a single step, similar to how humans approach challenging problems Wei et al.
+For step-wise reasoning, previous studies Wei et al.
 (2022); Zhao et al.
@@
 (2024); Yao et al.
-(2023).
-For retrieval-augmented generation (RAG), it is known that retrieving relevant external knowledge can mitigate hallucination and deal with timely knowledge, which is especially important for science domain Lewis et al.
+(2023) have shown that solving complex problems through a step-by-step process is more effective than attempting to solve them in a single step, mirroring how humans approach challenging problems.
+For retrieval-augmented generation (RAG), it is known that retrieving relevant external knowledge can mitigate hallucinations and help models adapt to newly emerging information, which is especially important in the scientific domain Lewis et al.
 (2020b); Bang et al.
@@
 (2024).
-To utilize the benefits of both worlds, recent works have adopted RAG for step-wise reasoning on various domains (Jeong et al., 2024; Zhang et al., 2025a).
-However, most of the works focus on relatively simple multi-hop question answering tasks or non-scientific domains (Guan et al., 2025; Jin et al., 2025; Wang et al., 2025) or assume task-related retrieval corpus which is hard to obtain compared to in-the-wild retrieval source such as Wikipedia (Zhong et al., 2025; Qian et al., 2023; Anand et al., 2023).
-Solving challenging scientific reasoning tasks such as graduate-level biology, physics, etc, using an in-the-wild corpus is very challenging since merely retrieving superficial knowledge is insufficient for these tasks.
-Instead, the retrieved information should contain relevant logics needed to solve the problem (Rueda et al., 2025).
+To utilize the benefits of both worlds, recent works have adopted RAG for step-wise reasoning across various applications Jeong et al.
+(2024); Zhang et al.
+(2025a).
+However, most studies focus on relatively simple multi-hop question answering tasks or non-scientific domains Guan et al.
+(2025); Jin et al.
+(2025); Wang et al.
+(2025), or assume the availability of task-specific retrieval corpora, which are more difficult to obtain than widely accessible web-based sources such as Wikipedia Zhong et al.
+(2025); Qian et al.
+(2023); Anand et al.
+(2023).
+Solving challenging scientific reasoning tasks,such as graduate-level biology or chemistry, using an in-the-wild corpus is difficult since merely retrieving superficial knowledge is insufficient.
+Instead, the retrieved information should contain relevant logical connections needed to solve the problem Rueda et al.
+(2025).
 Moreover, the knowledge required for each intermediate step can vary significantly even within the same problem.
@@
 Our framework is divided into three steps: (1) problem decomposition, (2) logical query generation, and (3) logical retrieval.
-For problem decomposition, LLMs decompose the original question into subquestions along with search queries that can assist in solving each reasoning step.
-Then, for each search query, LLMs convert the search query into a logical query, which is a query that contains logical explanations.
-Since logical queries contain more information about what external knowledge is needed to solve the problem, they enable retrievers to retrieve based on logic similarity rather than superficial domain similarity.
+For problem decomposition, LLMs break down the original question into subquestions along with search queries that can assist in solving each reasoning step.
+Then, for each search query, LLMs transform it into a logical query—one that includes the underlying logical explanations.
+Since logical queries contain more information about what external knowledge is needed to solve the problem, they enable retrieval based on logical rather than superficial domain similarity.
 Finally, we perform retrieval separately for each step, obtaining documents that are logically relevant to the particular step.
@@
 (2021); Wang et al.
-(2024b).RAISE consistently outperforms various baselines approaches that include applying RAG or problem decomposition.
+(2024b).
+Across these datasets, RAISE results in consistent performance improvement compared to various baseline approaches that include applying RAG or problem decomposition.
 This indicates that our framework effectively retrieves relevant documents that are needed to solve scientific reasoning tasks step-by-step.
 Through further analysis, we observe that RAISE retrieves more logically relevant documents for each step compared to baselines, highlighting that logical similarity is crucial for these challenging reasoning tasks in addition to domain similarity.
-Our code is available at https://github.com/minhaeoh/RAISE.
 2 Preliminary
@@
 $\text{sim}(x,d)=f_{q}(x)^{\top}f_{d}(d).$
-Top-$k$ documents are retrieved, and a generator (e.g., BART Lewis et al. (2020a), T5 Raffel et al. (2020)) produces $y$ based on both $x$ and $d_{i}$.
+[…]
+Retrieval in In-the-Wild Settings.
+We use the term “in-the-wild” to refer to open-domain corpora like Wikipedia that are not tailored for specific tasks or domains.
+Unlike curated corpora, they require retrieving logically relevant evidence from a large, diverse, and often tangential pool of content, making retrieval and reasoning more challenging.
 Figure 2: Examples comparing query generation methods (Step-Back, HyDE, and RAISE) for the same subquestion.
@@
 Rather than retrieving documents based on the entire problem as a single query, RAISE decomposes the problem into intermediate reasoning steps and performs targeted retrieval for each.
-The overview of RAISE is provided in Figure 1 and Algorithm 1.
+The overview of RAISE is provided in Figure 1 and Algorithm 1 in Appendix A.
 The method consists of three main stages:
@@
 Formally, we generate:
-$\{(r_{i},q_{i})\}_{i=1}^{n}\sim P_{\theta}(\cdot\mid x,p_{1})$ (1)
+$\{(r_{i},q_{i})\}_{i=1}^{n}\sim P_{\theta}(\cdot\mid x,p_{1}).$ (1)
 While recent work has increasingly explored problem decomposition, many retrieval-based approaches still operate at the problem level, retrieving information using a single query without explicitly modeling intermediate reasoning steps.
@@
 Instead, this step constructs a logically coherent query that aligns with the reasoning intent of the subquestion.
-The model is prompted with both $q_{i}$ and $r_{i}$, along with a reformulation prompt $p_{2}$ (e.g.
-add essential scientific or mathematical explanation the search result would contain):
+The model is prompted with both $q_{i}$ and $r_{i}$, along with a reformulation prompt $p_{2}$ (e.g., add essential scientific or mathematical explanation the search result would contain):
 $q_{i}^{*}\sim P_{\theta}(\cdot\mid q_{i},r_{i},p_{2})$ (2)
@@
 By combining both, we enable the model to synthesize a logical query that is purpose-driven and aligned with the knowledge required to answer $r_{i}$.
-Even if the reformulated query $q_{i}^{*}$ contains facutal inaccuracies, it tends to retrieve passages from corpus $\mathcal{C}$ that are logically relevant and supportive of the reasoning required for solving the original problem.
-Figure 2 presents example queries generated by RAISE ,Step-back and HyDE, illustrating RAISE’s ability to generate logical queries that are well-aglined with the reasoning intent.
-Logical Retrieval.    The reformulated query $q_{i}^{*}$ is used to retrieve external knowledge:
+Even if the reformulated query $q_{i}^{*}$ contains factual inaccuracies, it tends to retrieve passages from a corpus $\mathcal{C}$ that are logically relevant and supportive of the reasoning required for solving the original problem.
+Figure 2 presents example queries generated by RAISE, Step-Back+RAG, and HyDE, illustrating RAISE’s ability to generate logical queries that are well-aligned with the reasoning intent.
+Logical Retrieval.
+The reformulated query $q_{i}^{*}$ is used to retrieve external knowledge:
 $D_{i}=\mathcal{R}(q_{i}^{*},\mathcal{C},k)$ (3)
-Here, $\mathcal{R}$ is a top-$k$ retrieval function over in-the-wild corpus $\mathcal{C}$ (e.g., Wikipedia), and $D_{i}=\{d_{i1},\dots,d_{ik}\}$ denotes the retrieved documents.
-The goal is not to retrieve direct answers but to obtain passages that contain relevant background knowledge or reasoning procedures necessary to solve $r_{i}$.
-Also, since irrelevant documents might be also retrieved, we set a retrieval threshold $T$, where we retrieve documents that have higher retrieval similarity than $T$.
+[…]
 After retrieval, for each subquestion $r_{i}$, the model predicts its solution $a_{i}$.
@@
 This step-wise approach ensures that the model constructs its final answer based on grounded and explicitly structured intermediate reasoning.
-Table 1: Comparison of various reasoning strategies across GPQA, SuperGPQA, and MMLU. RAISE consistently outperforms other approaches for scientific reasoning benchmarks.
-Direct Reasoning Decomposed Reasoning
-CoT PS Step-back CoT+RAG Least-to-Most Step-back RAG RAISE
-GPQA
-Overall 42.42 39.39 43.94 45.96 44.95 44.44 45.95 51.01
-SuperGPQA
-science-hard 7.54 9.05 9.05 7.54 6.03 5.03 6.03 10.05
-science-middle 15.08 14.07 17.59 12.56 14.57 15.08 15.08 19.60
-engineering-hard 7.54 4.02 5.03 7.54 10.05 7.54 8.04 10.55
-MMLU
-(Pro) Chemistry 25.44 26.50 26.86 25.18 24.56 22.70 22.97 28.36
-(STEM) College Chemistry 49.50 45.00 44.00 43.00 45.40 43.00 46.00 51.00
+GPQA SuperGPQA MMLU
+Overall
+science-
+hard
+science-
+middle
+engineering-
+hard
+Chemistry
+Biology
+College Chemistry
+Direct
+CoT 42.42 4.52 15.08 6.53 25.44 51.88 49.50
+Direct+RAG
+CoT+RAG 45.96 7.54 12.56 7.54 25.18 54.39 43.00
+Decomposed
+Least-to-Most 44.95 6.03 14.57 10.05 24.56 53.97 45.40
+Step-Back 44.44 5.03 15.08 6.03 22.70 56.49 43.00
+Decomposed+RAG
+Least-to-Most+RAG 45.95 6.03 14.57 8.04 22.97 58.02 46.00
+Step-Back+RAG 43.43 5.53 15.58 9.05 23.06 56.34 43.00
+HyDE 46.46 7.54 13.07 7.04 22.97 57.88 49.00
+Ours
+RAISE
+51.01
+(+9.8%)
+10.05
+(+33.3%)
+19.60
+(+25.8%)
+10.55
+(+5.0%)
+28.36
+(+11.5%)
+59.27
+(+2.2%)
+51.00
+(+3.0%)
+Table 1: Comparison of various reasoning strategies across GPQA, SuperGPQA, and MMLU.
+The underscore indicates the highest score among baselines, while the boldface indicates the highest score overall.
+The value in parentheses below the RAISE score represents the performance gain compared to the best baseline.
+RAISE consistently outperforms other approaches for scientific reasoning benchmarks.
 4 Experiments
 4.1 Experimental Setup
-Datasets.
+4.1.1 Datasets.
 To evaluate our method in settings that require multi-step scientific reasoning, we select three datasets focused on high-level STEM and professional reasoning.
-We provide more details on the evaluation dataset in Appendix A.
+We provide more details on the datasets in Appendix B.
   • GPQA (Graduate-level Physics Question Answering) Rein et al.
@@
 We use this dataset as our primary benchmark.
-  • SuperGPQA (Team et al., 2025): This dataset evaluates the knowledge and reasoning abilities of Large Language Models across 285 graduate-level disciplines.
+  • SuperGPQA Team et al.
+(2025): This dataset evaluates the knowledge and reasoning abilities of LLMs across 285 graduate-level disciplines.
 We focus on three subsets that are challenging enough and require step-by-step reasoning in the science domain: science-hard, science-middle, and engineering-hard.
-  • MMLU (Hendrycks et al., 2021; Wang et al., 2024b): Among various domains and difficulty levels of MMLU (Hendrycks et al., 2021) and MMLU-Pro (Wang et al., 2024b), we focus on chemistry domain since it requires step-by-step reasoning with knowledge of specialized terminology.
-Retriever Configuration.
-We adopt Dense Passage Retrieval (DPR) Karpukhin et al.
-(2020) as our retriever, using the publicly available model trained on the Natural Questions (NQ) dataset Kwiatkowski et al.
-(2019).
-Language Model.
+  • MMLU Hendrycks et al.
+(2021); Wang et al.
+(2024b): We focus on a subset of science-related tasks in MMLU-Pro, specifically chemistry and biology, which are well-suited for evaluating step-by-step reasoning with specialized knowledge.
+Given the lower overall difficulty of MMLU-STEM compared to MMLU-Pro, we additionally include its chemistry category as a lightweight validation setting.
+4.1.2 Retriever Configuration.
+We adopt Dense Passage Retrieval (DPR) Karpukhin et al. (2020) as our retriever, using the publicly available model trained on the Natural Questions (NQ) dataset Kwiatkowski et al. (2019).
+4.1.3 Language Model.
 For GPQA, which is the primary focus of our study, we use Mistral Small 3.1-Instruct-2503 Mistral AI (2025), a 24B state-of-the-art small-scale open-sourced LLM that achieves strong performance on scientific reasoning tasks.
@@
 (2024) due to computational constraints, applying it to SuperGPQA, and MMLU.
-We provide detailed hyperparameters and experiment setups in Appendix B.
-4.2 Baselines
-To assess the importance of multi-step reasoning and step-aware retrieval, we conduct experiments with two groups of baselines: Direct Reasoning and Decomposed Reasoning.
-Direct Reasoning methods solve the problem in a single inference step without decomposing it into subquestions.
-This group includes three prompt-based approaches and one retrieval-augmented method.
-The three prompting methods-Chain-of-Thought (CoT) Wei et al.
-(2022), Plan-and-Solve (PS) Wang et al.
-(2023), Step-back Zheng et al.
-(2024) -differ in how the prompt guides the language models’ reasoning.
-Additionally, we include CoT+RAG Lewis et al.
-(2020b), a conventional retrieval-augmented method that retrieves documents using the original question as query.
-Decomposed Reasoning methods first break down the original problem into subquestions and solve them sequentially.
-This group includes Least-to-Most Zhou et al.
-(2023), which solves each subquestion directly without retreival, Step-back, which applies principle extraction at the subquestion level, and RAG, which retrieves and uses evidence for each subquestion individually.
-More detailed descriptions of each baseline can be found in Appendix B.1
+We provide detailed hyperparameters and experiment setups in Appendix C.
+4.1.4 Baselines.
+To assess the importance of multi-step reasoning and step-aware retrieval, we conduct experiments with four groups of baselines: Direct Reasoning, Direct Reasoning with RAG, Decomposed Reasoning, and Decomposed Reasoning with RAG.
+Direct Reasoning solves problems in a single step without decomposing them into subquestions, and includes Chain-of-Thought (CoT) Wei et al.
+(2022), which generates intermediate reasoning steps through prompting.
+A retrieval-augmented variant, Direct Reasoning with RAG, includes CoT+RAG Lewis et al.
+(2020b), which retrieves documents using the original question as the query.
+Decomposed Reasoning methods break down the original problem into subquestions and solve them sequentially.
+This includes Least-to-Most Zhou et al.
+(2023), which solves each subquestion directly, and Step-Back Zheng et al.
+(2024), which applies principle extraction at the subquestion level.
+A retrieval-augmented variant, Decomposed Reasoning with RAG, retrieves evidence for each subquestion and solves them step-by-step.
+This group includes Least-to-Most+RAG Lewis et al.
+(2020b), Step-Back+RAG Zheng et al.
+(2024), and HyDE Gao et al.
+(2023a), with the latter two enhancing retrieval relevance through query reformulation, making them strong baselines for comparison.
 4.2 Main Results
-As shown in Table 1, our proposed method, RAISE, consistently outperforms all baseline reasoning strategies across three benchmark datasets: GPQA, SuperGPQA, and MMLU.
-On GPQA, RAISE achieves 51.01%, surpassing the next-best method (CoT+RAG) by over 5 percentage points.
-On SuperGPQA, which consists of especially challenging science and engineering questions, RAISE achieves the highest performance across all subdomains, with gains especially notable in the science-middle subset (19.60%, +2 pts over the best baseline).
-On MMLU, RAISE also shows significant improvement, particularly in (Pro) Chemistry (28.36%) and STEM College Chemistry (51.00%), indicating its robustness across both professional and academic scientific reasoning tasks.
-This result shows that RAISE effectively decomposes scientific reasoning questions and retrieves documents that help the model to solve the questions.
-Unlike RAISE, decomposed reasoning methods do not always lead to better performance.
-This is often due to a lack of logically relevant knowledge especially for smaller LLMs, which limits their ability to reason effectively across multiple steps (Hosseini et al., 2024; Xu et al., 2025).
-In fact, the original papers of our baseline models that use decomposed reasoning mostly focused on proprietary large LLMs while we focus on open-sourced smaller LLMs.
-Despite well-structured decomposition into subquestions, a lack of logically relevant knowledge may cause the model to produce hallucinations or unsupported reasoning steps.
-Especially for smaller models that lack sufficient prior knowledge, problem decomposition can become counterproductive unless logically relevant documents are retrieved to support the reasoning process.
-Also, we observe that in some cases, models without retrieval outperform their retrieval-augmented counterparts.
-While retrieval is generally expected to enhance performance by providing additional information, these results reveal that naively applying retrieval can be ineffective or even harmful.
-This issue becomes particularly pronounced in settings involving open-domain or in-the-wild corpora, where the risk of retrieving noisy or irrelevant content is high.
+4.2.1 Baseline Comparison.
+As shown in Table 1, our proposed method, RAISE, consistently outperforms all baseline reasoning strategies across three benchmark datasets of varying difficulty: GPQA, SuperGPQA, and MMLU, achieving an average performance improvement of 13% over the best baseline scores.
+Unlike other baselines whose performance varies depending on the dataset’s difficulty or type, RAISE consistently demonstrates robust performance and outperforms them across different domains, types, and levels of difficulty.
+Specifically, on challenging GPQA, RAISE achieves a score of 51.01, outperforming the next-best method (HyDE) by 9.8%.
+To assess the effectiveness of our logical query generation, we compare RAISE with three RAG-based decomposed reasoning baselines that differ in how they construct retrieval queries.
+Least-to-Most+RAG uses the subquestion itself as the query, Step-Back+RAG abstracts a general principle from the subquestion, and HyDE generates a hypothetical answer to use as the retrieval query.
+RAISE consistently outperforms all baselines across benchmarks, demonstrating the advantage of generating logically grounded queries that better align with the reasoning required to solve each subquestion.
+These results confirm that RAISE’s queries go beyond retrieving documents that are merely domain-relevant or superficially similar, enabling access to knowledge that is logically aligned with the problem-solving process.
+Unlike RAISE, decomposed reasoning methods do not always yield better performance, particularly for smaller LLMs that lack logically relevant knowledge Hosseini et al.
+(2024); Xu et al.
+(2025).
+While prior work Zhou et al.
+(2023) using decomposition focused on proprietary large models, our study targets open-source smaller LLMs.
+Without sufficient background knowledge, even well-structured subquestions can lead to hallucinations or unsupported reasoning, making decomposition potentially counterproductive unless relevant documents are retrieved.
+We also observe cases where models without retrieval outperform their retrieval-augmented variants.
+For example, on the SuperGPQA engineering-hard subset, Least-to-Most+RAG (8.04) performs worse than Least-to-Most (10.05).
+Although retrieval is intended to improve performance by providing additional information, naive retrieval can introduce noise, especially in open-domain settings where irrelevant content is common.
 Conventional RAG methods often retrieve documents that are only superficially related or worse, focus on irrelevant variables or distractors rather than the core principles required for reasoning.
-These findings highlight the importance of retrieving logically relevant knowledge, and further underscore the effectiveness of our method.
-Table 2: Comparison of decomposed reasoning strategies with retrieval. RAISE consistently outperforms baselines across datasets.
-GPQA SuperGPQA MMLU
-RAG 45.95 9.72 24.67
-Step-back+RAG 43.43 10.05 33.03
-HyDE 28.28 9.22 35.99
-RAISE 51.01 13.57 39.68
-4.4 Additional Results
-Impact of Logical Query Generation.
-To evaluate the impact of logical query generation in retrieval-augmented reasoning, we conduct an experiment comparing our method, RAISE, with three RAG-based decomposed reasoning baselines: RAG, step-back+RAG (Zheng et al., 2024), and HyDE (Gao et al., 2023a).
-While all methods incorporate problem decomposition and retrieval, they differ in how the retrieval query is constructed.
-RAISE begins with a short key phrase known as the initial query, and expands it into a logical query that captures not only domain similarity but also the underlying reasoning intent of the subquestion.
-In contrast, the standard RAG baseline directly uses the subquestion itself as the retrieval query.
-The step-back+RAG approach attempts to abstract a general principle from the subquestion and uses that principle as the query.
-HyDE generates a synthetic passage in response to the subquestion and uses this generated text as the retrieval query, aiming to enrich the query context.
-As shown in Table 2, RAISE outperforms all other RAG-based baselines across all benchmarks.
-These results confirm that RAISE’s logical query generation goes beyond retrieving documents that are merely domain-relevant or superficially similar, enabling the retrieval of knowledge that is logically aligned with the reasoning required to solve the problem.
-We conduct further analysis in Section 4.5.
-For SuperGPQA and MMLU, Table 2 reports average scores across subsets; full per-subset results are provided in the Appendix 5.
-Figure 3: Performance comparison between RAISE-Direct and RAISE across datasets.
-Table 3: Evaluation on GPQA with various LLMs with different scales: LLaMA 3.1-8B, GPT-4o mini, and Mistral Small 3.1. RAISE consistently shows the best performance across all settings.
-LLaMA GPT Mistral
-CoT 22.22 40.91 42.42
-PS 27.78 38.90 39.39
-Step-back 28.28 40.90 43.94
-CoT+RAG 23.23 40.40 45.96
-Least-to-Most 26.26 45.45 44.95
-Step-back (Decomp.) 28.28 42.42 44.44
-RAG 24.24 42.93 45.95
-RAISE 30.30 47.98 51.01
-Impact of Problem Decomposition.
-To further assess the importance of problem decomposition, we also evaluate a variant of our method that omits this step and directly performs logical query generation and retrieval without breaking the problem into subquestions, as shown in Figure 3.
-This version, referred to as RAISE-Direct, showed lower performance compared to the full version of RAISE.
-These results indicate that problem decomposition plays a critical role in guiding the retrieval process and structuring the reasoning pathway.
-Applying RAISE to various LLMs.
-To assess the generalizability of RAISE across different LLM scales, we evaluate its performance on GPQA using LLaMA 3.1-8B (Grattafiori et al., 2024) and GPT-4o mini (OpenAI, 2024), in addition to Mistral(used in our main experiments).
+These results highlight the importance of retrieving logically relevant knowledge and further support the effectiveness of our approach.
+4.2.2 Applying RAISE to various LLMs.
+To assess the generalizability of RAISE across different LLM scales, we evaluate its performance on GPQA using LLaMA 3.1-8B Grattafiori et al.
+(2024) and GPT-4o mini OpenAI (2024), in addition to Mistral (used in our main experiments).
 As shown in Table 2, RAISE demonstrates consistent improvements over other baselines, exhibiting a similar trend to our main results with Mistral-24B.
 This shows that the effect of RAISE is not limited to a specific type of LLM, but can be applied to various LLMs with different scales.
-Table 4: Examples where RAISE-retrieved documents provide logically relevant information for scientific reasoning compared to baseline RAG retrieval.
+LLaMA GPT Mistral
+Direct
+CoT 22.22 40.91 42.42
+Direct+RAG
+CoT+RAG 23.23 40.40 45.96
+Decomposed
+Least-to-Most 26.26 45.45 44.95
+Step-Back 28.28 42.42 44.44
+Decomposed+RAG
+Least-to-Most+RAG 24.24 42.93 45.95
+Step-Back+RAG 21.72 42.42 43.43
+HyDE 25.75 38.89 46.46
+Ours
+RAISE
+30.30
+(+7.1%)
+47.98
+(+5.3%)
+51.01
+(+9.8%)
+Table 2: Evaluation on GPQA with various LLMs with different scales: LLaMA 3.1-8B, GPT-4o mini, and Mistral Small 3.1.
+Underscore marks the best baseline; bold indicates the best overall.
+Values in parentheses under RAISE show gains over the top baseline.
+RAISE consistently shows the best performance across all settings.
 Subquestion RAG RAISE Explanation
@@
 The RAISE-retrieved document clearly identifies barium hydroxide as $\mathrm{Ba(OH)_{2}}$ and explains its dissociation behavior in water, directly supporting the calculation of $[\mathrm{OH}^{-}]$ concentration.
-How does the scalar potential (V) change when considering the retarded time $t_{r}$?
-Compensating variation in economics … additional money needed after a price change.
-The calculation of the retarded time $t_{r}=t-|\vec{r}-\vec{r^{\prime}}|/c$ is a speed-distance-time relation.
-The field observed at time $t$ must have been emitted at this earlier time.
-The RAISE-retrieved document directly explains the concept of retarded time in electrodynamics as the time delay due to the finite propagation speed of electromagnetic fields, which is essential for deriving the scalar potential $V(\vec{r},t)$ of a moving charge.
+Table 3: Examples where RAISE-retrieved documents provide logically relevant information for scientific reasoning compared to baseline RAG retrieval.
 4.3 Analysis of RAISE
+4.3.1 Impact of Problem Decomposition.
+To further assess the importance of problem decomposition, we also evaluate a variant of our method that omits this step and directly performs logical query generation and retrieval without breaking the problem into subquestions, as shown in Figure 3.
+This version, referred to as RAISE-Direct, showed lower performance compared to the full version of RAISE.
+These results indicate that problem decomposition plays a critical role in guiding the retrieval process and structuring the reasoning pathway.
+This suggests that for complex reasoning problems, decomposing the question and retrieving logical knowledge tailored to each subquestion is more effective than retrieving once based on the original question alone.
+This is likely because different reasoning steps often require distinct pieces of information that may not be jointly retrievable from a single query.
+Figure 3: Performance comparison between RAISE-Direct and RAISE across datasets.
+4.3.2 Qualitative Evaluation of Retrieved Documents
+We qualitatively demonstrate the examples when RAISE retrieves logically relevant documents compared to convential RAG in Table 3.
+While RAG often retrieves documents that are topically related yet fail to address the reasoning needs of the subquestion, RAISE consistently identifies documents that include essential scientific principles, mechanisms, or equations.
+For instance, in questions involving chemical reactions, RAISE surfaces materials that explain the specific reactivity or the retarded time calculation, whereas RAG may return vague definitions or unrelated economic concepts.
+These cases illustrate how RAISE’s retrieval is not only domain-aware but also aligned with the logical demands of solving complex scientific problems.
+Figure 4: Logical Relevancy of Retrieved Documents.
+Unlike other baselines, RAISE has higher ratio of documents that are logically relevant and lower ratio of documents that are irrelevant or superficially relevant.
 To further investigate our hypothesis that RAISE retrieves documents that are logically more relevant compared to other baselines, we use LLM-as-a-judge (GPT-4o-mini) to evaluate the logical relevancy of the retrieved documents.
@@
 This indicates that RAISE avoids retrieving documents that may interrupt the reasoning process for scientific reasoning through logical query generation.
-We also qualitatively demonstrate the examples when RAISE retrieves logically relevant documents compared to convential RAG in Table 4.
-While RAG often retrieves documents that are topically related yet fail to address the reasoning needs of the subquestion, RAISE consistently identifies documents that include essential scientific principles, mechanisms, or equations.
-For instance, in questions involving chemical reactions or physical formulas, RAISE surfaces materials that explain the specific reactivity or the retarded time calculation, whereas RAG may return vague definitions or unrelated economic concepts.
-These cases illustrate how RAISE’s retrieval is not only domain-aware but also aligned with the logical demands of solving complex scientific problems.
-Figure 4: Logical Relevancy of Retrieved Documents.
-Unlike other baselines, RAISE has higher ratio of documents that are logically relevant and lower ratio of documents that are irrelevant or superficially relevant.
+Additionally, since LLM-as-judge may struggle with assessing logical relevance in expert-level science problems like GPQA and SuperGPQA, we conducted human evaluations with 2–3 experts per question, including chemistry professors and PhDs.The results can be found in Appendix D.2.
+These human evaluation results further confirm that RAISE significantly outperforms other baselines in selecting logically relevant documents.
 5 Related Works
@@
 Unlike other domains, scientific reasoning requires not only step-by-step thinking, but also knowledge of specialized terminology and adaptation to continually evolving knowledge.
-Due to this challenging nature, many benchmarks have been proposed recently to tackle scientific reasoning with LLMs (Rein et al., 2024; Zhong et al., 2025; Zhang et al., 2025b; Wang et al., 2024a; Jiang et al., 2024).
-Many works enhance scientific reasoning capabilities of LLMs through domain-specific training (Taylor et al., 2022; Prabhakar et al., 2025; Zhang et al., 2024b), step-by-step reasoning (Rueda et al., 2025; Rein et al., 2024), or retrieval of external knowledge or tools (Ma et al., 2024; Zhong et al., 2025; Wellawatte et al., 2025; Li et al., 2025).
+Due to this challenging nature, many benchmarks have been proposed recently to tackle scientific reasoning with LLMs Rein et al.
+(2024); Zhong et al.
+(2025); Zhang et al.
+(2025b); Wang et al.
+(2024a); Jiang et al.
+(2024).
+Many works enhance scientific reasoning capabilities of LLMs through domain-specific training Taylor et al.
+(2022); Prabhakar et al.
+(2025); Zhang et al.
+(2024b), step-by-step reasoning Rueda et al.
+(2025); Rein et al.
+(2024), or retrieval of external knowledge or tools Ma et al.
+(2024); Zhong et al.
+(2025); Wellawatte et al.
+(2025); Li et al.
+(2025).
 Unlike previous works, we focus on applying step-by-step document retrieval from in-the-wild corpus without assuming access to well-curated and domain-specific corpus.
@@
 We investigate how external evidence can be integrated at each step to improve reasoning beyond what internal knowledge alone can achieve.
-Retrieval-Augmented Generation.
+Retrieval Augmented Generation.
 Retrieval-Augmented Generation (RAG) Lewis et al.
@@
 This enables the model to supplement its limited internal knowledge with external sources, leading to more complete problem solving, especially in complex, multi-step tasks.
-6 Conclusion
+6 Conclusion and Limitations
 We introduce RAISE, a step-by-step retrieval framework for scientific reasoning.
@@
 definition of specificalized terminology) but also logically relevant documents for each subquestion, assisting the step-by-step reasoning process required for scientific reasoning.
-Limitations
-In our study, we primarily conducted experiments using relatively smaller language models, mainly due to computational resource constraints.
-As a result, the performance observed may not fully reflect the potential of larger-scale models.
-Future work could explore scaling up the model size to assess its impact on reasoning capabilities and overall performance.
-For the retrieval process, we used only DPR as the retriever.
-While DPR has been widely recognized for its effectiveness, we believe that more recent retrieval models Gao et al.
+However, RAISE has some limitations and areas for future improvement.
+First, we used only DPR for retrieval.
+Incorporating recent models Gao et al.
 (2023b); Basu et al.
 (2024); Shi et al.
-(2023), could lead to a substantial improvement in performance.
-Incorporating these advanced retrievers remains a promising direction for future research.
-Furthermore, our work is particularly focused on scientific reasoning tasks.
-While this domain presents unique challenges and complexities, extending our approach to other domains such as legal, commonsense, or medical reasoning could be an interesting and valuable avenue for further investigation.
-Such cross-domain studies may help validate the generalizability and robustness of RAISE.
-Additionally, RAISE relies on multiple calls to large language models during the logical query generation phase.
-Although this multi-step querying is essential for generating effective logical queries that maximize retrieval efficiency, exploring strategies to reduce the number of LLM invocations, such as more efficient prompting, intermediate caching, or lightweight planning modules, could enhance the overall efficiency and practicality of our method in real-world applications.
-Appendix A Dataset Details
-GPQA Rein et al.
+(2023) may further improve performance.
+In addition, our study focuses on scientific reasoning.
+Applying RAISE to other domains such as legal, commonsense, or medical reasoning could help evaluate its generalizability.
+Moreover, RAISE requires multiple LLM calls during logical query generation.
+Reducing this cost through better prompting, caching, or lightweight planning could improve scalability and real-world usability.
+Appendix A RAISE Algorithm
+Input: Original question $x$, prompts $\mathcal{P}=\{p_{1},p_{2},p_{3},p_{4}\}$, corpus $\mathcal{C}$
+Output: Final answer $y$
+Step 1: Problem Decomposition
+Generate subquestions and initial queries:
+$\{(r_{i},q_{i})\}_{i=1}^{n}\sim P_{\theta}(\cdot\mid x,p_{1})$
+for $i=1$ to $n$ do
+ Step 2: Logical Query Generation
+   Reformulate initial query:
+    $q_{i}^{*}\sim P_{\theta}(\cdot\mid r_{i},q_{i},p_{2})$
+ Step 3: Knowledge Retrieval
+[…]
+    $D_{i}=\mathcal{R}(q_{i}^{*},\mathcal{C},k)$
+ Step 4: Subquestion Answering
+ if $i=1$ then
+       $a_{i}\sim P_{\theta}(\cdot\mid x,r_{1},D_{1},p_{3})$
+ else
+       $a_{i}\sim P_{\theta}(\cdot\mid x,\{(r_{j},a_{j})\}_{j=1}^{i-1},r_{i},D_{i},p_{3})$
+   end if
+end for
+Step 5: Final Answer Composition
+Generate final answer using all subanswers:
+$y\sim P_{\theta}(\cdot\mid x,\{(r_{i},a_{i})\}_{i=1}^{n},p_{4})$
+Algorithm 1 RAISE Inference Procedure
+Appendix B Dataset Details
+GPQA
+Rein et al.
 (2024) This dataset consists of physics, biology, and chemistry questions written by domain experts.
@@
 To ensure consistency during evaluation, we proprocessed each question by randomly shuffling the correct answer along with the three distractors, and assigning them uniformly to choice labels (A) through (D).
-SuperGPQA (Team et al., 2025)    SuperGPQA is a large-scale benchmark designed to evaluate graduate-level reasoning across 13 disciplines, 72 fields, and 285 graduate-level disciplines.
+SuperGPQA
+Team et al.
+(2025) SuperGPQA is a large-scale benchmark designed to evaluate graduate-level reasoning across 13 disciplines, 72 fields, and 285 graduate-level disciplines.
 In alignment with the scientific reasoning focus of our work, we select science and engineering domains for evaluation.
@@
 Specifically, our experiments include 199 examples each from science-hard, science-middle, and engineering-hard subsets.
-MMLU (Hendrycks et al., 2021; Wang et al., 2024b)    The MMLU benchmark covers a wide range of subjects across multiple domains.
+MMLU
+Hendrycks et al.
+(2021); Wang et al.
+(2024b) The MMLU benchmark covers a wide range of subjects across multiple domains.
 For out experiments, we focus on STEM and Professional categories.
 The STEM contains university-level science and engineering subjects such as college mathematics and computer science, while the Professional category covers specialized fields that typically require professional training or advanced education, including law, medicine, and chemistry.
-We specifically select two subsets: college chemistry from MMLU-STEM and professional chemistry from MMLU-Pro.
+We specifically select three subsets: college chemistry from MMLU-STEM and professional chemistry and biology from MMLU-Pro.
 These subsets are chosen to evaluate our method’s ability to perform scientific reasoning in both academic and professional contexts involving complex domain knowledge.
-Appendix B Experiment Details
-B.1 Baseline Details
+Appendix C Experiment Details
+C.1 Baseline Details
 CoT Wei et al.
@@
 (2023) We apply Chain-of-Thought prompting for direct reasoning, where the model is encouraged to explicitly generate intermediate reasoning steps through prompting (Think step by step).
-PS Wang et al. (2023) Plan-and-Solve prompting is a direct reasoning strategy where the model first outlines a plan and then executes it step by step. We use the same prompt as the original paper.
-Step-back (Direct reasoning) Zheng et al.
-(2024) Step-back prompting guides reasoning using abstracted principles derived from the original question.
-In our implementation, we follow the original Step-back paper’s prompting strategy to extract high-level principles necessary for solving the question.
-These extracted principles are then provided together with the original question as input to guide the model’s reasoning process.
-CoT+RAG Lewis et al.
+CoT+RAG
+Lewis et al.
 (2020b) We implement CoT+RAG by combining Chain-of-Thought prompting with retrieval, where the model is prompted to solve the problem step-by-step while also leveraging external knowledge.
 Specifically, we provide the model with a CoT-style prompt encouraging step-by-step reasoning, alongside the original question and documents retrieved using the original question as the search query.
-Least-to-Most Zhou et al.
+Least-to-Most
+Zhou et al.
 (2023) Least-to-Most is a decomposed reasoning strategy that breaks down a complex problem into a sequence of simpler subquestions, which are then solved sequentially without retrieval augmentation.
 This subquestion decomposition pipeline serves as the foundational structure for other decomposed reasoning methods as well.
-Step-back (Decomposed reasoning) Zheng et al.
+Step-Back
+(Decomposed reasoning) Zheng et al.
 (2024) We implement Step-Back for decomposed reasoning by applying the Step-Back prompting method to each subquestion in a decomposed reasoning framework.
@@
 Specifically, after decomposing the original question into subquestions, we use the Step-Back prompting strategy to extract a high-level principle for each subquestion, and then provide the subquestion along with its corresponding principle to guide the model’s reasoning.
-RAG (Decomposed reasoning) Liu et al.
+Least-to-Most+RAG
+(Decomposed reasoning with RAG) Liu et al.
 (2025) We implement RAG by first decomposing the original problem into subquestions and then retrieving documents using each subquestion as a query.
 The retrieved documents are provided to the model along with the corresponding subquestion to support its reasoning.
-B.2 Retriever Configuration
+Step-Back+RAG
+(Decomposed reasoning with RAG) Zheng et al.
+(2024) We extend the Step-Back prompting strategy to a retrieval-augmented setting for fair comparison with our method.
+After decomposing the original question into subquestions, we generate a principle abstraction for each subquestion using Step-Back prompting, and use it as a query to retrieve evidence.
+The retrieved documents are then provided alongside the original subquestion to guide the model’s reasoning.
+HyDE
+(Decomposed reasoning with RAG) Gao et al.
+(2023a) We apply the HyDE approach to each subquestion in a decomposed reasoning framework.
+For each subquestion, the model first generates a hypothetical answer, which is then used as a query to retrieve supporting documents.
+The retrieved evidence, together with the subquestion, is provided to the model to support step-by-step reasoning.
+C.2 Retriever Configuration
 We use the pre-trained DPR encoder from the ’facebook/dpr-question_encoder-single-nq-base’ model Karpukhin et al.
@@
 Throughout all experiment, we retrieve top-10 documents per query.
-B.3 RAISE
-To reduce the impact of potentially irrelevant documents by DPR, we apply a similarity threshold $T$ in RAISE.Specifically, we discard any retrieved passage whose DPR similarity score falls below $T$.
+To reduce the impact of potentially irrelevant documents by DPR, we apply a similarity threshold $T$ in RAISE.
+Specifically, we discard any retrieved passage whose DPR similarity score falls below $T$.
 DPR similarity is computed as the inner product between L2-normalized query and passage embeddings.
@@
 For MMLU-STEM (college chemistry), we use a slightly lower threshold of $T=0.80$, considering that the questions are generally simpler than those in other datasets.
-B.4 Algorithm
-Input: Original question $x$, prompts $\mathcal{P}=\{p_{1},p_{2},p_{3},p_{4}\}$, corpus $\mathcal{C}$
-Output: Final answer $y$
-Step 1: Problem Decomposition
-Generate subquestions and initial queries:
-$\{(r_{i},q_{i})\}_{i=1}^{n}\sim P_{\theta}(\cdot\mid x,p_{1})$
-for $i=1$ to $n$ do
-       Step 2: Logical Query Generation
-       Reformulate initial query:
-        $q_{i}^{*}\sim P_{\theta}(\cdot\mid r_{i},q_{i},p_{2})$
-      Step 3: Knowledge Retrieval
-       Retrieve top-$k$ documents:
-        $D_{i}=\mathcal{R}(q_{i}^{*},\mathcal{C},k)$
-      Step 4: Subquestion Answering
-       if $i=1$ then
-              $a_{i}\sim P_{\theta}(\cdot\mid x,r_{1},D_{1},p_{3})$
-      else
-              $a_{i}\sim P_{\theta}(\cdot\mid x,\{(r_{j},a_{j})\}_{j=1}^{i-1},r_{i},D_{i},p_{3})$
-       end if
-end for
-Step 5: Final Answer Composition
-Generate final answer using all subanswers:
-$y\sim P_{\theta}(\cdot\mid x,\{(r_{i},a_{i})\}_{i=1}^{n},p_{4})$
-Algorithm 1 RAISE Inference Procedure
-B.5 Prompts
-B.5.1 Baseline Prompts
-Prompt for CoT You are solving a multiple choice question.
-Think step by step and show your reasoning clearly.
-At the end, state your answer in the format: "The final answer is (X)".
-Here, X must be the correct letter choice.
-Question: [Problem here] Answer Choices: [Answer choices here] Solution:
-Prompt for Plan-and-Solve Let’s first understand the problem and devise a plan to solve the problem.
-Then, let’s carry out the plan and solve the problem step by step.
-Use the following format exactly: <Plan> Step 1: [state what you will do] Step 2: [state next reasoning step] … (as needed) <Solution> Step 1: [carry out Step 1] Step 2: [carry out Step 2] … (as needed) "The final answer is (X)." Here, X must be the correct letter choice.
-Question: [Problem here] Answer Choices: [Answer choices here] Solution:
-Prompt for Step-back principle abstraction You are an expert at Science.
-You are given a Science problem.
-Your task is to extract the Science concepts and principles involved in solving the problem.
-What are the principles behind this question?
-End your response with "End of generation" after you answer the instructions.
-Question: [Subquestion here] Principles Involved:
-Prompt for Step-back You are an expert at Science.
-You are given a Science problem and a set of principles involved in solving the problem.
-Solve the problem step by step by folowing the principles.
-At the end, state your answer in the format: "The final answer is (X)".
-Here, X must be the correct letter choice.
-Question: [Problem here] Principles: [Principles here] Answer Choices: [Answer choices here] Solution:
-Prompt for hyDE query generation Generate a paragraph that answers the question. End your response with "End of generation" after you answer the instructions. Question: [Subquestion here] Explanation:
-B.5.2 RAISE Prompts
-Prompt for Problem Decomposition You are given a multiple-choice question.
-Break this problem into essential subquestions that directly help solve the original problem.
-Each subquestion MUST also include its search query.
-Each search query should reflect scientific or mathematical knowledge needed to answer the subquestion.
-STRICT FORMAT REQUIREMENTS: 1.
-For each subquestion, you MUST provide exactly two parts in this order: - The subquestion - A search query for that subquestion 2.
-Use EXACTLY this format for each subquestion: Subquestion 1: [your specific subquestion] Search Query for Subquestion 1: [Write a search query someone might realistically use to learn how to answer this subquestion] Question: [Problem here] Answer Choices: [Answer choices here]
-Prompt for Logical Query Generation You are given a subquestion and a search query.
-The search query is a realistic phrase that someone might use to find knowledge or reasoning support to answer the subquestion.
-Your task is to anticipate what essential scientific or mathematical explanation the search result would contain, and write it concisely (2–3 sentences).
-Focus only on the core concept or principle that would help answer the subquestion.
-Avoid restating the subquestion, and do not include unrelated or overly general information.
-Subquestion: [Subquestion resulting from Problem Decomposition] Search Query: [Search query resulting from Problem Decomposition] Explanation:
-Prompt for Solving Subquestions with Documents You are solving a multiple-choice question.
-Question is decomposed into several subquestions.
-You will be given: 1.
-The original multiple-choice question 2.
-Previous subquestions and their solutions (if any) 3.
-The current subquestion to solve 4.
-Documents that are relevant to the current subquestion Your task: - Carefully read the original question, any previous subquestion and their solutions, and the current subquestion - Use the information from the retrieved documents to solve the current subquestion - Also use your existing knowledge to solve the current subquestion - Your solution should be detailed and logically structured Documents: [Retrieved document] Question: [Problem here] Answer Choices: [Answer choices here] Previous subquestions and their solutions: [Previously generated subquestions and solutions] Current subquestion to solve: Subquestion [Step num]: [Subquestion] Subquestion [Step num] Solution:
-Prompt for Generating Final Answer You are solving a multiple-choice question.
-Question is decomposed into several subquestions.
-Each subquestion has already been solved.
-Your task is to carefully read the original question and the several subquestion solutions, then use them to determine the final answer.
-Think step by step and then finish your answer with "The final answer is (X)" where X is the correct letter choice.
-Original Question: Question: [Problem here] Answer Choices: [Answer choices here] Subquestions and Solutions: [Generated stepwise subproblems and solutions] Final Solution:
-Prompt for Evaluation with GPT You are given the following three items: - Original Problem: [Problem here] - Subquestion: [Subquestion here] - Retrieved Document: [Document here] Your task is to evaluate how helpful the retrieved document is for answering the subquestion.
-Please follow these instructions: - Do not just check if the topic is related.
-- Instead, check if the document includes information that helps someone reason through and solve the subquestion.
-- Focus on whether the document supports actual thinking or steps needed to get the answer.
-Give your final judgment using only one of the following ratings: - "No relevance at all" - does not have any domain similarity - "Superficially relevant" - has domain similarity (only superficially) but does not have any logical relevance to the subquestion.
-For example, the document might mention the same topic as the subquestion, but it does not provide any information that helps solve the subquestion.
-- "Partially relevant" - has domain similarity and has some logical relevance to the subquestion.
-For example, the document might provide some information that helps solve the subquestion, but it does not provide all the logical steps needed.
-- "Fully relevant" - has domain similarity and has almost all logical relevance to the subquestion.
-For example, the document provides enough relevant logical steps to solve the subquestion.
-Then explain your reasoning briefly.
-Output Format: Helpfulness Rating: <one of the 4 options above> Explanation: <your short explanation>
-Appendix C Additional Results
+Appendix D Additional Results
+D.1 Further Analysis on GPQA
 Figure 5: Domain-wise accuracy comparison between CoT and RAISE on the GPQA Diamond subset
@@
 These results demonstrate RAISE’s robustness and its ability to generalize across diverse areas of graduate-level scientific reasoning.
-Table 5 provides the full subset results for each benchmark, complementing the averaged scores shown in Table 2.
-Across all datasets, RAISE consistently outperforms other retrieval-augmented baselines.
-The consistent improvement across subsets supports our central hypothesis that constructing queries enriched with logical intent is essential for retrieving logically relevant knoweldge.
-Table 5: Comparison of decomposed reasoning strategies with retrieval. Subset-level results for each benchmark, complementing the averaged scores reported in the main text.
-RAG Step-back +RAG HyDE RAISE
-GPQA
-Overall 45.95 43.43 48.48 51.01
-SuperGPQA
-science-hard 6.03 5.53 7.54 10.55
-science-middle 15.08 15.58 13.07 19.60
-engineering-hard 8.04 9.05 7.04 10.55
-MMLU
-(Pro) Chemistry 22.53 23.06 22.97 28.36
-(STEM) College Physics 46.00 43.00 49.00 51.00
+D.2 Human Evaluation of Retrieved Documents
+Figure 6: Human Evaluation of the Logical Relevance of Retrieved Documents Aligned with the results from the LLM-as-a-judge evaluation of logical relevancy, RAISE shows a higher proportion of logically relevant documents and a lower proportion of irrelevant or superficially relevant ones.
+Previously, in Section 4.3.2, we evaluated the logical relevancy of documents retrieved by the baseline methods and RAISE using the LLM-as-a-judge approach.
+However, since our domain includes complex, expert-level questions, and LLM-based evaluations may overlook domain-specific reasoning and often rely on surface-level features, we supplemented our analysis with a small-scale human evaluation of 20 subquestion–document pairs.
+Each pair was assessed by at least three annotators, including Ph.D.
+students and a faculty member in chemistry, with the method provenance concealed to maintain objectivity.
+The results are presented in Figure 6.
+As also discussed in the LLM-as-a-judge results, the human evaluation indicates that RAISE produces significantly fewer irrelevant documents compared to all other methods, while achieving the highest proportion of logically relevant documents.
+Although limited in scale due to time and cost constraints, we believe this evaluation provides meaningful human validation of RAISE’s effectiveness and serves as a valuable complement to the LLM-based assessments.
+Appendix E Prompts
+E.1 Baseline Prompts
+You are solving a multiple choice question. Think step by step and show your reasoning clearly.
+At the end, state your answer in the format: "The final answer is (X)".
+Here, X must be the correct letter choice.
+Question: [Problem here]
+Answer Choices: [Answer choices here]
+Solution:
+Figure 7: Prompt for CoT
+You are an expert at Science. You are given a Science problem.
+Your task is to extract the Science concepts and principles involved in solving the problem.
+What are the principles behind this question?
+End your response with "End of generation" after you answer the instructions.
+Question: [Subquestion here]
+Principles Involved:
+Figure 8: Prompt for Step-Back Principle Abstraction
+You are an expert at Science. You are given a Science problem and a set of principles involved in solving the problem.
+Solve the problem step by step by following the principles.
+At the end, state your answer in the format: "The final answer is (X)".
+Here, X must be the correct letter choice.
+Question: [Problem here]
+Principles: [Principles here]
+Answer Choices: [Answer choices here]
+Solution:
+Figure 9: Prompt for Step-Back
+Generate a paragraph that answers the question.
+End your response with "End of generation" after you answer the instructions.
+Question: [Subquestion here]
+Explanation:
+Figure 10: Prompt for HyDE Query Generation
+E.2 RAISE Prompts
+You are given a multiple-choice question.
+Break this problem into essential subquestions that directly help solve the original problem.
+Each subquestion MUST also include its search query.
+Each search query should reflect scientific or mathematical knowledge needed to answer the subquestion.
+STRICT FORMAT REQUIREMENTS:
+1. For each subquestion, you MUST provide exactly two parts in this order:
+- The subquestion
+- A search query for that subquestion
+2. Use EXACTLY this format for each subquestion:
+Subquestion 1: [your specific subquestion]
+Search Query for Subquestion 1: [Write a search query someone might realistically use to learn how to answer this subquestion]
+Question: [Problem here]
+Answer Choices: [Answer choices here]
+Figure 11: Prompt for Problem Decomposition
+You are given a subquestion and a search query.
+The search query is a realistic phrase that someone might use to find knowledge or reasoning support to answer the subquestion.
+Your task is to anticipate what essential scientific or mathematical explanation the search result would contain, and write it concisely (2–3 sentences).
+Focus only on the core concept or principle that would help answer the subquestion.
+Avoid restating the subquestion, and do not include unrelated or overly general information.
+Subquestion: [Subquestion resulting from Problem Decomposition]
+Search Query: [Search query resulting from Problem Decomposition]
+Explanation:
+Figure 12: Prompt for Logical Query Generation
+You are solving a multiple-choice question. The question is decomposed into several subquestions. You will be given:
+ 1. The original multiple-choice question
+ 2. Previous subquestions and their solutions (if any)
+ 3. The current subquestion to solve
+ 4. Documents that are relevant to the current subquestion
+Your task:
+- Carefully read the original question, any previous subquestions and their solutions, and the current subquestion.
+- Use the information from the retrieved documents to solve the current subquestion.
+- Also use your existing knowledge to solve the current subquestion.
+- Your solution should be detailed and logically structured.
+Documents: [Retrieved document]
+Question: [Problem here]
+Answer Choices: [Answer choices here]
+Previous subquestions and their solutions:
+[Previously generated subquestions and solutions]
+Current subquestion to solve:
+Subquestion [Step num]: [Subquestion]
+Subquestion [Step num] Solution:
+Figure 13: Prompt for Solving Subquestions with Documents
+You are solving a multiple-choice question.
+The question is decomposed into several subquestions.
+Each subquestion has already been solved.
+Your task is to carefully read the original question and the several subquestion solutions, then use them to determine the final answer.
+Think step by step and then finish your answer with "The final answer is (X)" where X is the correct letter choice.
+Original Question:
+Question: [Problem here]
+Answer Choices: [Answer choices here]
+Subquestions and Solutions:
+[Generated stepwise subproblems and solutions]
+Final Solution:
+Figure 14: Prompt for Generating Final Answer
+You are given the following three items:
+- Original Problem: [Problem here]
+- Subquestion: [Subquestion here]
+- Retrieved Document: [Document here]
+Your task is to evaluate how helpful the retrieved document is for answering the subquestion.
+Please follow these instructions:
+- Do not just check if the topic is related.
+- Instead, check if the document includes information that helps someone reason through and solve the subquestion.
+- Focus on whether the document supports actual thinking or steps needed to get the answer.
+Give your final judgment using only one of the following ratings:
+- "No relevance at all" – does not have any domain similarity
+- "Superficially relevant" – has domain similarity (only superficially) but does not have any logical relevance to the subquestion.
+For example, the document might mention the same topic as the subquestion, but it does not provide any information that helps solve the subquestion.
+- "Partially relevant" – has domain similarity and has some logical relevance to the subquestion.
+For example, the document might provide some information that helps solve the subquestion, but it does not provide all the logical steps needed.
+- "Fully relevant" – has domain similarity and has almost all logical relevance to the subquestion. For example, the document provides enough relevant logical steps to solve the subquestion.
+Then explain your reasoning briefly.
+Output Format:
+Helpfulness Rating: <one of the 4 options above>
+Explanation: <your short explanation>
+Figure 15: Prompt for Evaluation with GPT

### v3

@@
 Scientific reasoning requires not only long-chain reasoning processes, but also knowledge of domain-specific terminologies and adaptation to updated findings.
-To address these challenges in scientific reasoning, we introduce RAISE, a step-by-step retrieval-augmented framework which retrieves logically relevant documents from an in-the-wild corpus.
+To deal with these challenges for scientific reasoning, we introduce RAISE, a step-by-step retrieval-augmented framework which retrieves logically relevant documents from in-the-wild corpus.
 RAISE is divided into three steps: problem decomposition, logical query generation, and logical retrieval.
@@
 We analyze that unlike other baselines, RAISE retrieves documents that are not only similar in terms of the domain knowledge, but also documents logically more relevant.
-RAISE: Enhancing Scientific Reasoning in LLMs via
-Step-by-Step Retrieval
-Minhae Oh1, Jeonghye Kim2, Nakyung Lee1, Donggeon Seo3, Taeuk Kim3, Jungwoo Lee1, 1Seoul National University, 2KAIST, 3Hanyang University
+Codes are available at https://github.com/minhaeoh/RAISE.
 1 Introduction
-Recently, large language models (LLMs) have demonstrated remarkable capabilities in scientific domains that require intricate reasoning and the integration of diverse knowledge sources Zhang et al.
-(2024c, a); Prabhakar et al.
-(2025); Rueda et al.
-(2025).
-In contrast to many other domains, scientific reasoning is especially demanding due to the need for advanced reasoning skills, knowledge of specialized terminology, and awareness of continually evolving knowledge.
-To enhance the capability of LLMs to solve scientific reasoning problems, two approaches are widely used: (1) step-wise reasoning, and (2) retrieval-augmented generation.
-For step-wise reasoning, previous studies Wei et al.
-(2022); Zhao et al.
-(2024a); Jin et al.
-(2024); Lai et al.
-(2024); Yao et al.
-(2023) have shown that solving complex problems through a step-by-step process is more effective than attempting to solve them in a single step, mirroring how humans approach challenging problems.
-For retrieval-augmented generation (RAG), it is known that retrieving relevant external knowledge can mitigate hallucinations and help models adapt to newly emerging information, which is especially important in the scientific domain Lewis et al.
-(2020b); Bang et al.
-(2025); Zhong et al.
-(2025); Xiong et al.
-(2024).
-To utilize the benefits of both worlds, recent works have adopted RAG for step-wise reasoning across various applications Jeong et al.
-(2024); Zhang et al.
-(2025a).
-However, most studies focus on relatively simple multi-hop question answering tasks or non-scientific domains Guan et al.
-(2025); Jin et al.
-(2025); Wang et al.
-(2025), or assume the availability of task-specific retrieval corpora, which are more difficult to obtain than widely accessible web-based sources such as Wikipedia Zhong et al.
-(2025); Qian et al.
-(2023); Anand et al.
-(2023).
+Large language models (LLMs) have shown strong potential for scientific reasoning, which demands advanced reasoning skills, domain-specific terminology, and up-to-date knowledge [45, 42, 27, 30].
+Two common strategies are step-wise reasoning, which solves complex problems through structured intermediate steps [37, 47, 12, 18, 41], and retrieval-augmented generation (RAG), which mitigates hallucinations by providing external evidence [21, 1, 50, 39].
+Recent work combines them, but often targets simpler multi-hop QA or assumes curated, task-specific corpora [9, 44, 5, 11, 34], unlike open-domain sources such as Wikipedia.
 Solving challenging scientific reasoning tasks, such as graduate-level biology or chemistry, using an in-the-wild corpus is difficult since merely retrieving superficial knowledge is insufficient.
-Instead, the retrieved information should contain relevant logical connections needed to solve the problem Rueda et al.
-(2025).
+Instead, the retrieved information should contain relevant logical connections needed to solve the problem [30].
 Moreover, the knowledge required for each intermediate step can vary significantly even within the same problem.
@@
 The question of what to search for and how to retrieve the appropriate external knowledge for each step when solving scientific reasoning tasks is underexplored.
-Figure 1: Overview of RAISE.
-RAISE is divided into three steps: (1) Problem Decomposition, (2) Logical Query Generation, and (3) Logical Retrieval.
-Given an scientific reasoning question, we decompose the problem into multiple subquestions and search queries.
-Then, we generate a logical query that contains logic-enriched contexts.
-Finally, we retrieve relevant documents from in-the-wild corpus for each subquestion and condition the documents to assist the step-by-step reasoning for scientific reasoning problems.
+Figure 1: Overview of RAISE. RAISE is divided into three steps: (1) Problem Decomposition, (2) Logical Query Generation, and (3) Logical Retrieval.
 To address these challenges, we introduce RAISE (Step-by-Step Retrieval-Augmented Inference for Scientific rEasoning), a retrieval-augmented framework tailored for step-wise scientific reasoning.
-Our framework is divided into three steps: (1) problem decomposition, (2) logical query generation, and (3) logical retrieval.
-For problem decomposition, LLMs break down the original question into subquestions along with search queries that can assist in solving each reasoning step.
-Then, for each search query, LLMs transform it into a logical query—one that includes the underlying logical explanations.
-Since logical queries contain more information about what external knowledge is needed to solve the problem, they enable retrieval based on logical rather than superficial domain similarity.
-Finally, we perform retrieval separately for each step, obtaining documents that are logically relevant to the particular step.
+Our framework consists of three stages: (1) problem decomposition, where LLMs break down the original question into subquestions along with search queries; (2) logical query generation, which reformulates each search query into a logic-enriched query that captures the reasoning needed to solve the subquestion; and (3) logical retrieval, which retrieves step-specific documents from an open-domain corpus, ensuring the retrieved information is logically relevant rather than superficially domain similar.
 Instead of assuming task-relevant or well-curated retrieval source, such as question-answer pool of relevant domains, we retrieve from in-the-wild source such as Wikipedia, which enables applying to challenging real-world scenarios.
-We evaluate our framework on three challenging scientific reasoning benchmarks: GPQA Rein et al.
-(2024), SuperGPQA Team et al.
-(2025), and MMLU Hendrycks et al.
-(2021); Wang et al.
-(2024b).
-Across these datasets, RAISE results in consistent performance improvement compared to various baseline approaches that include applying RAG or problem decomposition.
-This indicates that our framework effectively retrieves relevant documents that are needed to solve scientific reasoning tasks step-by-step.
-Through further analysis, we observe that RAISE retrieves more logically relevant documents for each step compared to baselines, highlighting that logical similarity is crucial for these challenging reasoning tasks in addition to domain similarity.
+Evaluated on GPQA, SuperGPQA, and MMLU, RAISE consistently outperforms baselines using either RAG or problem decomposition alone, demonstrating its ability to retrieve step-specific, logically relevant information essential for solving complex scientific reasoning tasks.
+While this work focuses on scientific reasoning benchmarks, similar challenges arise in mathematical reasoning, which also requires precise, multi-step logical inference.
 2 Preliminary
@@
 $\text{sim}(x,d)=f_{q}(x)^{\top}f_{d}(d).$
-[…]
+Top-$k$ documents are retrieved, and a generator (e.g., BART [20], T5 [28]) produces $y$ based on both $x$ and $d_{i}$.
 Retrieval in In-the-Wild Settings.
@@
 Unlike curated corpora, they require retrieving logically relevant evidence from a large, diverse, and often tangential pool of content, making retrieval and reasoning more challenging.
-Figure 2: Examples comparing query generation methods (Step-Back, HyDE, and RAISE) for the same subquestion.
 3 RAISE
 We propose RAISE (Step-by-Step Retrieval-Augmented Inference for Scientific rEasoning), a retrieval-augmented generation framework for scientific reasoning designed to support multi-step reasoning through fine-grained, step-aware retrieval.
-Rather than retrieving documents based on the entire problem as a single query, RAISE decomposes the problem into intermediate reasoning steps and performs targeted retrieval for each.
+The method consists of three main stages: (1) Problem Decomposition, (2) Logical Query Generation, (3) Logical Retreival.
 The overview of RAISE is provided in Figure 1 and Algorithm 1 in Appendix A.
-The method consists of three main stages:
-  1. (1)
-    Problem Decomposition, where the original question $x$ is broken down into subquestions $r_{1},\dots,r_{n}$ along with corresponding search queries $q_{1},\dots,q_{n}$ that are intended to retrieve relevant knowledge for each step;
-  2. (2)
-    Logical Query Generation, where each search query $q_{i}$ is further enriched into a logic aware explanation $q_{i}^{*}$, which serves as the final retrieval query; and
-  3. (3)
-    Logical Retrieval, where external knowledge $D_{i}$ is retrieved for each subquestion $r_{i}$ and used to generate the subanswer $a_{i}$.
 Problem Decomposition.
-RAISE begins with the decomposition of a scientific reasoning problem $x$ into a sequence of intermediate subquestions $\{r_{i}\}_{i=1}^{n}$, each corresponding to a distinct reasoning step.
-This planning process is performed by prompting a language model $P_{\theta}$ with $x$ and prompt $p_{1}$ to identify the logical structure required to solve the problem.
-Formally, we generate:
-$\{(r_{i},q_{i})\}_{i=1}^{n}\sim P_{\theta}(\cdot\mid x,p_{1}).$ (1)
-While recent work has increasingly explored problem decomposition, many retrieval-based approaches still operate at the problem level, retrieving information using a single query without explicitly modeling intermediate reasoning steps.
-In contrast, this stage produces a structured sequence of subquestions that reflects the multi-step nature of scientific reasoning.
-Along with each subquestion $r_{i}$, a corresponding search query $q_{i}$ is generated to guide retrieval.
+RAISE decomposes the problem $x$ into subquestions $r_{1},\dots,r_{n}$ with corresponding search queries $q_{1},\dots,q_{n}$, forming a structured sequence for step-wise retrieval, unlike conventional single-query approaches.
 These queries are not used directly for retrieval but rather serve as an initial query for the next stage.
@@
 In the second stage, each initial search query $q_{i}$ and its corresponding subquestion $r_{i}$ are jointly used to generate a logically enriched logical query $q_{i}^{*}$.
-Search queries $q_{i}$ are short key phrases, where the logical process is absent.
-Therefore, using $q_{i}$ for retrieval will focus on domain similarity instead of logical similarity.
-Instead, this step constructs a logically coherent query that aligns with the reasoning intent of the subquestion.
-The model is prompted with both $q_{i}$ and $r_{i}$, along with a reformulation prompt $p_{2}$ (e.g., add essential scientific or mathematical explanation the search result would contain):
-$q_{i}^{*}\sim P_{\theta}(\cdot\mid q_{i},r_{i},p_{2})$ (2)
-Using only the subquestion $r_{i}$ to generate a query often results in noisy or overly specific formulations that include problem-specific variables or conditions irrelevant to general knowledge retrieval.
-On the other hand, using only the initial query $q_{i}$ lacks sufficient context to capture the intended reasoning path.
-By combining both, we enable the model to synthesize a logical query that is purpose-driven and aligned with the knowledge required to answer $r_{i}$.
+Since initial queries $q_{i}$ lacks reasoning context and subquestions $r_{i}$ alone can be noisy or overly specific, neither $q_{i}$ nor $r_{i}$ alone is sufficient for effective retrieval.
+By combining both, we generate logical queries that better capture the reasoning intent and retrieve logically relevant knowledge for solving each step.
+The model is prompted with both $q_{i}$ and $r_{i}$, along with a reformulation prompt $p_{2}$.
 Even if the reformulated query $q_{i}^{*}$ contains factual inaccuracies, it tends to retrieve passages from a corpus $\mathcal{C}$ that are logically relevant and supportive of the reasoning required for solving the original problem.
-Figure 2 presents example queries generated by RAISE, Step-Back+RAG, and HyDE, illustrating RAISE’s ability to generate logical queries that are well-aligned with the reasoning intent.
+Figure 4 in Appendix C.2 presents example queries generated by RAISE, Step-Back+RAG, and HyDE, illustrating RAISE’s ability to generate logical queries that are well-aligned with the reasoning intent.
 Logical Retrieval.
-The reformulated query $q_{i}^{*}$ is used to retrieve external knowledge:
-$D_{i}=\mathcal{R}(q_{i}^{*},\mathcal{C},k)$ (3)
-[…]
-After retrieval, for each subquestion $r_{i}$, the model predicts its solution $a_{i}$.
-Importantly, the prediction is conditioned not only on the retrieved documents $D_{i}$, but also on the original question $x$ and the previously answered subquestions $\{(r_{j},a_{j})\}_{j=1}^{i-1}$.
-We write:
-$a_{i}\sim P_{\theta}(\cdot\mid x,\{(r_{j},a_{j})\}_{j=1}^{i-1},r_{i},D_{i},p_{3})$ (4)
-Finally, all subanswers are aggregated to predict the final answer $y$:
-$y\sim P_{\theta}(\cdot\mid x,\{(r_{i},a_{i})\}_{i=1}^{n},p_{4})$ (5)
-This step-wise approach ensures that the model constructs its final answer based on grounded and explicitly structured intermediate reasoning.
+External knowledge $D_{i}$ is retrieved for each subquestion $r_{i}$ from in-the-wild corpus $\mathcal{C}$ (e.g., Wikipedia) and used to generate the subanswer $a_{i}$.
+We retrieve background knowledge for each subquestion using a similarity threshold $T$ to filter irrelevant documents.
+After retrieval, for each subquestion $r_{i}$, the model predicts its solution $a_{i}$ using $D_{i}$, the original question $x$, and the previous steps.
+Finally, all subanswers are combined to generate the final answer $y$.
 GPQA SuperGPQA MMLU
-Overall
-science-
-hard
-science-
-middle
-engineering-
-hard
-Chemistry
-Biology
-College Chemistry
+Overall science- hard science- middle engineering- hard (Pro) Chemistry (Pro) Biology (STEM) College Chemistry
 Direct
@@
 Ours
-RAISE
-51.01
-(+9.8%)
-10.05
-(+33.3%)
-19.60
-(+25.8%)
-10.55
-(+5.0%)
-28.36
-(+11.5%)
-59.27
-(+2.2%)
-51.00
-(+3.0%)
+RAISE 51.01 (+9.8%) 10.05 (+33.3%) 19.60 (+25.8%) 10.55 (+5.0%) 28.36 (+11.5%) 59.27 (+2.2%) 51.00 (+3.0%)
 Table 1: Comparison of various reasoning strategies across GPQA, SuperGPQA, and MMLU.
-The underscore indicates the highest score among baselines, while the boldface indicates the highest score overall.
-The value in parentheses below the RAISE score represents the performance gain compared to the best baseline.
+The underscore marks the best baseline, boldface the best overall, and parentheses show RAISE’s gain over the top baseline.
 RAISE consistently outperforms other approaches for scientific reasoning benchmarks.
-4 Experiments
+4 Experiment
 4.1 Experimental Setup
-4.1.1 Datasets.
-To evaluate our method in settings that require multi-step scientific reasoning, we select three datasets focused on high-level STEM and professional reasoning.
-We provide more details on the datasets in Appendix B.
-  • GPQA (Graduate-level Physics Question Answering) Rein et al.
-(2024): A collection of graduate-level physics, biology, and chemistry questions requiring multi-step reasoning and access to external scientific fact.
-We use this dataset as our primary benchmark.
-  • SuperGPQA Team et al.
-(2025): This dataset evaluates the knowledge and reasoning abilities of LLMs across 285 graduate-level disciplines.
-We focus on three subsets that are challenging enough and require step-by-step reasoning in the science domain: science-hard, science-middle, and engineering-hard.
-  • MMLU Hendrycks et al.
-(2021); Wang et al.
-(2024b): We focus on a subset of science-related tasks in MMLU-Pro, specifically chemistry and biology, which are well-suited for evaluating step-by-step reasoning with specialized knowledge.
-Given the lower overall difficulty of MMLU-STEM compared to MMLU-Pro, we additionally include its chemistry category as a lightweight validation setting.
-4.1.2 Retriever Configuration.
-We adopt Dense Passage Retrieval (DPR) Karpukhin et al. (2020) as our retriever, using the publicly available model trained on the Natural Questions (NQ) dataset Kwiatkowski et al. (2019).
-4.1.3 Language Model.
-For GPQA, which is the primary focus of our study, we use Mistral Small 3.1-Instruct-2503 Mistral AI (2025), a 24B state-of-the-art small-scale open-sourced LLM that achieves strong performance on scientific reasoning tasks.
-For the remaining tasks, we use the smaller LLaMA 3.1-8B model Grattafiori et al.
-(2024) due to computational constraints, applying it to SuperGPQA, and MMLU.
-We provide detailed hyperparameters and experiment setups in Appendix C.
-4.1.4 Baselines.
-To assess the importance of multi-step reasoning and step-aware retrieval, we conduct experiments with four groups of baselines: Direct Reasoning, Direct Reasoning with RAG, Decomposed Reasoning, and Decomposed Reasoning with RAG.
-Direct Reasoning solves problems in a single step without decomposing them into subquestions, and includes Chain-of-Thought (CoT) Wei et al.
-(2022), which generates intermediate reasoning steps through prompting.
-A retrieval-augmented variant, Direct Reasoning with RAG, includes CoT+RAG Lewis et al.
-(2020b), which retrieves documents using the original question as the query.
-Decomposed Reasoning methods break down the original problem into subquestions and solve them sequentially.
-This includes Least-to-Most Zhou et al.
-(2023), which solves each subquestion directly, and Step-Back Zheng et al.
-(2024), which applies principle extraction at the subquestion level.
-A retrieval-augmented variant, Decomposed Reasoning with RAG, retrieves evidence for each subquestion and solves them step-by-step.
-This group includes Least-to-Most+RAG Lewis et al.
-(2020b), Step-Back+RAG Zheng et al.
-(2024), and HyDE Gao et al.
-(2023a), with the latter two enhancing retrieval relevance through query reformulation, making them strong baselines for comparison.
+Datasets.
+We evaluate on three scientific benchmarks: GPQA, SuperGPQA, MMLU, which cover graduate-level STEM and professional science tasks that require multi-step scientific reasoning.
+Retriever and Language Models.
+We adopt Dense Passage Retrieval (DPR) [14] trained on the Natural Questions (NQ) dataset [17] as our retriever.
+For GPQA, our primary benchmark, we use Mistral Small 3.1-Instruct-2503 [25] (24B), while for SuperGPQA and MMLU we use the lighter LLaMA 3.1-8B model [4] due to computational limits.
+Baselines.
+To assess the importance of multi-step reasoning and step-aware retrieval, we conduct experiments with four groups of baselines: Direct Reasoning(CoT [37]), Direct Reasoning with RAG(CoT+RAG [21]), Decomposed Reasoning(Least-to-Most [51], Step-Back [49]), and Decomposed Reasoning with RAG.
+The last group retrieves evidence for each subquestion and solves them step-by-step.
+This group includes Least-to-Most+RAG [21], Step-Back+RAG, and HyDE [3], with the latter two improving retrieval relevance through query reformulation, making them strong baselines.
+Further details about datasets, retriever, model settings, and baselines are provided in Appendix B.
 4.2 Main Results
-4.2.1 Baseline Comparison.
 As shown in Table 1, our proposed method, RAISE, consistently outperforms all baseline reasoning strategies across three benchmark datasets of varying difficulty: GPQA, SuperGPQA, and MMLU, achieving an average performance improvement of 13% over the best baseline scores.
 Unlike other baselines whose performance varies depending on the dataset’s difficulty or type, RAISE consistently demonstrates robust performance and outperforms them across different domains, types, and levels of difficulty.
-Specifically, on challenging GPQA, RAISE achieves a score of 51.01, outperforming the next-best method (HyDE) by 9.8%.
+Furthermore, we confirm that these improvements hold across models of different scales, including smaller LLaMA-8B and GPT-4o mini, as shown in Appendix C.1, demonstrating that RAISE’s effectiveness is not tied to a specific LLM architecture or size.
 To assess the effectiveness of our logical query generation, we compare RAISE with three RAG-based decomposed reasoning baselines that differ in how they construct retrieval queries.
@@
 These results confirm that RAISE’s queries go beyond retrieving documents that are merely domain-relevant or superficially similar, enabling access to knowledge that is logically aligned with the problem-solving process.
-Unlike RAISE, decomposed reasoning methods do not always yield better performance, particularly for smaller LLMs that lack logically relevant knowledge Hosseini et al.
-(2024); Xu et al.
-(2025).
-While prior work Zhou et al.
-(2023) using decomposition focused on proprietary large models, our study targets open-source smaller LLMs.
-Without sufficient background knowledge, even well-structured subquestions can lead to hallucinations or unsupported reasoning, making decomposition potentially counterproductive unless relevant documents are retrieved.
-We also observe cases where models without retrieval outperform their retrieval-augmented variants.
-For example, on the SuperGPQA engineering-hard subset, Least-to-Most+RAG (8.04) performs worse than Least-to-Most (10.05).
-Although retrieval is intended to improve performance by providing additional information, naive retrieval can introduce noise, especially in open-domain settings where irrelevant content is common.
-Conventional RAG methods often retrieve documents that are only superficially related or worse, focus on irrelevant variables or distractors rather than the core principles required for reasoning.
-These results highlight the importance of retrieving logically relevant knowledge and further support the effectiveness of our approach.
-4.2.2 Applying RAISE to various LLMs.
-To assess the generalizability of RAISE across different LLM scales, we evaluate its performance on GPQA using LLaMA 3.1-8B Grattafiori et al.
-(2024) and GPT-4o mini OpenAI (2024), in addition to Mistral (used in our main experiments).
+Qualitative examples further support this finding, as shown in Appendix C.2, where RAISE retrieves passages containing essential scientific mechanisms while conventional RAG often returns vague or unrelated content.
+Unlike RAISE, decomposed reasoning methods do not always yield better performance, particularly for smaller open-source LLMs that lack sufficient background knowledge [7, 40].
+While decomposition can help structure reasoning, without access to relevant external knowledge, smaller models may produce hallucinations or unsupported intermediate steps, sometimes leading to worse performance than direct reasoning.
+Moreover, even when retrieval is added, naive RAG can introduce additional noise.
+In such cases, the retriever may surface superficially related or distracting content rather than the core principles needed for reasoning, which can ultimately harm performance.
+This highlights the importance of retrieving logically relevant knowledge rather than merely domain-related content, a challenge that RAISE directly addresses.
+5 Analysis of RAISE
+Figure 2: Performance comparison between RAISE-Direct and RAISE across datasets.
+To further assess the importance of problem decomposition, we also evaluate a variant of our method that omits this step and directly performs logical query generation and retrieval without breaking the problem into subquestions, as shown in Figure 2.
+This version, referred to as RAISE-Direct, showed lower performance compared to the full version of RAISE.
+These results indicate that problem decomposition plays a critical role in guiding the retrieval process and structuring the reasoning pathway.
+This suggests that for complex reasoning problems, decomposing the question and retrieving logical knowledge tailored to each subquestion is more effective than retrieving once based on the original question alone.
+This is likely because different reasoning steps often require distinct pieces of information that may not be jointly retrievable from a single query.
+We also analyze the quality of the retrieved documents.
+Using both an LLM-as-a-judge and a small-scale human evaluation, we find that RAISE consistently retrieves fewer irrelevant or superficial documents and more passages that directly support reasoning (Appendix C.3).
+These results confirm that RAISE’s gains stem from retrieving logically aligned knowledge rather than merely domain-related content.
+6 Conclusion
+We introduce RAISE, a step-by-step retrieval framework for scientific reasoning.
+We first decompose the problem into multiple subquestions and search queries, and then generate logical queries and retrieve logically relevant documents from in-the-wild corpus.
+We demonstrate the effectiveness of RAISE on three scientific reasoning benchmarks by comparing with various baselines.
+Our analysis shows that RAISE retrieves documents that are not only relevant in terms of the domain (e.g.
+definition of specificalized terminology) but also logically relevant documents for each subquestion, assisting the step-by-step reasoning process required for scientific reasoning.
+Although our experiments focus on scientific reasoning, the stepwise logical retrieval in RAISE is broadly applicable to other domains such as mathematical problem solving, which also demands precise multi-step inference.
+Acknowledgments
+This work is in part supported by the National Research Foundation of Korea (NRF, RS-2024-00451435(20%), RS-2024-00413957(20%)), Institute of Information & communications Technology Planning & Evaluation (IITP, RS-2021-II212068(10%), RS-2025-02305453(15%), RS-2025-02273157(15%), RS-2025-25442149(10%) RS-2021-II211343(10%)) grant funded by the Ministry of Science and ICT (MSIT), Institute of New Media and Communications(INMAC), and the BK21 FOUR program of the Education, Artificial Intelligence Graduate School Program (Seoul National University), and Research Program for Future ICT Pioneers, Seoul National University in 2025.
+Appendix A RAISE Algorithm
+Input: Original question $x$, prompts $\mathcal{P}=\{p_{1},p_{2},p_{3},p_{4}\}$, corpus $\mathcal{C}$
+Output: Final answer $y$
+Step 1: Problem Decomposition
+Generate subquestions and initial queries:
+$\{(r_{i},q_{i})\}_{i=1}^{n}\sim P_{\theta}(\cdot\mid x,p_{1})$
+for $i=1$ to $n$ do
+ Step 2: Logical Query Generation
+   Reformulate initial query:
+    $q_{i}^{*}\sim P_{\theta}(\cdot\mid r_{i},q_{i},p_{2})$
+ Step 3: Knowledge Retrieval
+   Retrieve top-$k$ documents:
+    $D_{i}=\mathcal{R}(q_{i}^{*},\mathcal{C},k)$
+ Step 4: Subquestion Answering
+ if $i=1$ then
+       $a_{i}\sim P_{\theta}(\cdot\mid x,r_{1},D_{1},p_{3})$
+ else
+       $a_{i}\sim P_{\theta}(\cdot\mid x,\{(r_{j},a_{j})\}_{j=1}^{i-1},r_{i},D_{i},p_{3})$
+   end if
+end for
+Step 5: Final Answer Composition
+Generate final answer using all subanswers:
+$y\sim P_{\theta}(\cdot\mid x,\{(r_{i},a_{i})\}_{i=1}^{n},p_{4})$
+Algorithm 1 RAISE Inference Procedure
+Appendix B Experiment Details
+B.1 Dataset Details
+GPQA
+[29] This dataset consists of physics, biology, and chemistry questions written by domain experts.
+We use GPQA diamond subset, which consist of 198 high-quality questions selected based on human performance.
+Specifically, this subset includes questions that both experts answer correctly while the majority of non-experts fail to solve.
+Each question typically demands multi-step reasoning, precise formula manipulation, and access to external scientific facts (e.g., physical constants, definitions).
+Due to its alignment with our target setting, GPQA serves as the primary evaluation benchmark throughout our experiments.
+For GPQA, the original dataset does not include standardized multiple-choice labeled as (A), (B), (C) and (D).
+To ensure consistency during evaluation, we proprocessed each question by randomly shuffling the correct answer along with the three distractors, and assigning them uniformly to choice labels (A) through (D).
+SuperGPQA
+[32] SuperGPQA is a large-scale benchmark designed to evaluate graduate-level reasoning across 13 disciplines, 72 fields, and 285 graduate-level disciplines.
+In alignment with the scientific reasoning focus of our work, we select science and engineering domains for evaluation.
+Each domain is further divided by three difficulty levels(easy, medium, and hard).
+To reduce computational overhead while maintaining consistency, we randomly sample 199 questions per subset using a fixed seed (42).
+Specifically, our experiments include 199 examples each from science-hard, science-middle, and engineering-hard subsets.
+MMLU
+[6, 36] The MMLU benchmark covers a wide range of subjects across multiple domains.
+For out experiments, we focus on STEM and Professional categories.
+The STEM contains university-level science and engineering subjects such as college mathematics and computer science, while the Professional category covers specialized fields that typically require professional training or advanced education, including law, medicine, and chemistry.
+We specifically select three subsets: college chemistry from MMLU-STEM and professional chemistry and biology from MMLU-Pro.
+These subsets are chosen to evaluate our method’s ability to perform scientific reasoning in both academic and professional contexts involving complex domain knowledge.
+B.2 Baseline Details
+CoT [37, 16] We apply Chain-of-Thought prompting for direct reasoning, where the model is encouraged to explicitly generate intermediate reasoning steps through prompting (Think step by step).
+CoT+RAG
+[21] We implement CoT+RAG by combining Chain-of-Thought prompting with retrieval, where the model is prompted to solve the problem step-by-step while also leveraging external knowledge.
+Specifically, we provide the model with a CoT-style prompt encouraging step-by-step reasoning, alongside the original question and documents retrieved using the original question as the search query.
+Least-to-Most
+[51] Least-to-Most is a decomposed reasoning strategy that breaks down a complex problem into a sequence of simpler subquestions, which are then solved sequentially without retrieval augmentation.
+This subquestion decomposition pipeline serves as the foundational structure for other decomposed reasoning methods as well.
+Step-Back
+(Decomposed reasoning) [49] We implement Step-Back for decomposed reasoning by applying the Step-Back prompting method to each subquestion in a decomposed reasoning framework.
+While the original Step-Back paper does not cover the application of this method to decomposed subquestions, we extend it for a fair comparison with our approach.
+Specifically, after decomposing the original question into subquestions, we use the Step-Back prompting strategy to extract a high-level principle for each subquestion, and then provide the subquestion along with its corresponding principle to guide the model’s reasoning.
+Least-to-Most+RAG
+(Decomposed reasoning with RAG) [23] We implement RAG by first decomposing the original problem into subquestions and then retrieving documents using each subquestion as a query.
+The retrieved documents are provided to the model along with the corresponding subquestion to support its reasoning.
+Step-Back+RAG
+(Decomposed reasoning with RAG) [49] We extend the Step-Back prompting strategy to a retrieval-augmented setting for fair comparison with our method.
+After decomposing the original question into subquestions, we generate a principle abstraction for each subquestion using Step-Back prompting, and use it as a query to retrieve evidence.
+The retrieved documents are then provided alongside the original subquestion to guide the model’s reasoning.
+HyDE
+(Decomposed reasoning with RAG) [3] We apply the HyDE approach to each subquestion in a decomposed reasoning framework.
+For each subquestion, the model first generates a hypothetical answer, which is then used as a query to retrieve supporting documents.
+The retrieved evidence, together with the subquestion, is provided to the model to support step-by-step reasoning.
+B.3 Retriever Configuration
+We use the pre-trained DPR encoder from the ’facebook/dpr-question_encoder-single-nq-base’ model [14], which is a BERT-based encoder trained for open-domain question answering.
+This encoder is trained on the Natural Question (NQ) dataset [17] and is designed to map questions into 768-dimensional dense vector representations for retrieval.
+For the retrieval corpus, we use the preprocessed Wikipedia passages provided by ’facebook/wiki_dpr’ [14], a corpus widely used to evaluate DPR-based retrieval models.
+This corpus is constructed from the December 20, 2018 Wikipedia dump, where each article is split into multiple, disjoint text blocks of 100 words, resulting in approximately 21 million passages.
+Each passage is accompanied by the title of the wikipedia page it comes from along with DPR embedding.
+To enable efficient retrieval over the passage embeddings, we use an exact FAISS index.
+FAISS (Facebook AI Similarity Search) [13, 2] is widely used library for fast similarity search over dense vectors.
+Throughout all experiment, we retrieve top-10 documents per query.
+To reduce the impact of potentially irrelevant documents by DPR, we apply a similarity threshold $T$ in RAISE.
+Specifically, we discard any retrieved passage whose DPR similarity score falls below $T$.
+DPR similarity is computed as the inner product between L2-normalized query and passage embeddings.
+Higher scores indicate greater semantic similarity, with values closer to 1 representing stronger alignment between the query and passage.
+We set $T=0.84$ for GPQA, SuperGPQA, and MMLU-Pro, which are composed of more challenging reasoning problems.
+For MMLU-STEM (college chemistry), we use a slightly lower threshold of $T=0.80$, considering that the questions are generally simpler than those in other datasets.
+Appendix C Additional Results
+C.1 Applying RAISE to various LLMs.
+To assess the generalizability of RAISE across different LLM scales, we evaluate its performance on GPQA using LLaMA 3.1-8B [4] and GPT-4o mini [26], in addition to Mistral (used in our main experiments).
 As shown in Table 2, RAISE demonstrates consistent improvements over other baselines, exhibiting a similar trend to our main results with Mistral-24B.
@@
 Ours
-RAISE
-30.30
-(+7.1%)
-47.98
-(+5.3%)
-51.01
-(+9.8%)
+RAISE 30.30 (+7.1%) 47.98 (+5.3%) 51.01 (+9.8%)
 Table 2: Evaluation on GPQA with various LLMs with different scales: LLaMA 3.1-8B, GPT-4o mini, and Mistral Small 3.1.
@@
 RAISE consistently shows the best performance across all settings.
+C.2 Qualitative Evaluation of Retrieved Documents
 Subquestion RAG RAISE Explanation
@@
 Barium hydroxide can be prepared by dissolving BaO in water… The Ba centers adopt a square anti-prismatic geometry.
-The RAISE-retrieved document clearly identifies barium hydroxide as $\mathrm{Ba(OH)_{2}}$ and explains its dissociation behavior in water, directly supporting the calculation of $[\mathrm{OH}^{-}]$ concentration.
-Table 3: Examples where RAISE-retrieved documents provide logically relevant information for scientific reasoning compared to baseline RAG retrieval.
-4.3 Analysis of RAISE
-4.3.1 Impact of Problem Decomposition.
-To further assess the importance of problem decomposition, we also evaluate a variant of our method that omits this step and directly performs logical query generation and retrieval without breaking the problem into subquestions, as shown in Figure 3.
-This version, referred to as RAISE-Direct, showed lower performance compared to the full version of RAISE.
-These results indicate that problem decomposition plays a critical role in guiding the retrieval process and structuring the reasoning pathway.
-This suggests that for complex reasoning problems, decomposing the question and retrieving logical knowledge tailored to each subquestion is more effective than retrieving once based on the original question alone.
-This is likely because different reasoning steps often require distinct pieces of information that may not be jointly retrievable from a single query.
-Figure 3: Performance comparison between RAISE-Direct and RAISE across datasets.
-4.3.2 Qualitative Evaluation of Retrieved Documents
-We qualitatively demonstrate the examples when RAISE retrieves logically relevant documents compared to convential RAG in Table 3.
+The RAISE-retrieved document clearly identifies barium hydroxide as $\mathrm{Ba(OH)_{2}}$ and explains its dissociation behavior in water, directly supporting the calculation of $\mathrm{OH}^{-}]$ concentration.
+Figure 3: Examples where RAISE-retrieved documents provide logically relevant information for scientific reasoning compared to baseline RAG retrieval.
+We qualitatively demonstrate the examples when RAISE retrieves logically relevant documents compared to convential RAG in Figure 3.
 While RAG often retrieves documents that are topically related yet fail to address the reasoning needs of the subquestion, RAISE consistently identifies documents that include essential scientific principles, mechanisms, or equations.
@@
 These cases illustrate how RAISE’s retrieval is not only domain-aware but also aligned with the logical demands of solving complex scientific problems.
+Figure 4: Examples comparing query generation methods (Step-Back+RAG, HyDE, and RAISE) for the same subquestion.
+Both Step-Back+RAG and HyDE are methods that reformulate the original query to retrieve more relevant documents.
+These methods are included as baselines in the main comparison table.
 Figure 5: Logical Relevancy of Retrieved Documents.
 Unlike other baselines, RAISE has higher ratio of documents that are logically relevant and lower ratio of documents that are irrelevant or superficially relevant.
+Figure 6: Human Evaluation of the Logical Relevance of Retrieved Documents Aligned with the results from the LLM-as-a-judge evaluation of logical relevancy, RAISE shows a higher proportion of logically relevant documents and a lower proportion of irrelevant or superficially relevant ones.
+C.3 Evaluation of Logical Relevancy of Retrieved Documents
 To further investigate our hypothesis that RAISE retrieves documents that are logically more relevant compared to other baselines, we use LLM-as-a-judge (GPT-4o-mini) to evaluate the logical relevancy of the retrieved documents.
@@
 This indicates that RAISE avoids retrieving documents that may interrupt the reasoning process for scientific reasoning through logical query generation.
-Additionally, since LLM-as-judge may struggle with assessing logical relevance in expert-level science problems like GPQA and SuperGPQA, we conducted human evaluations with 2–3 experts per question, including chemistry professors and PhDs.The results can be found in Appendix D.2.
-These human evaluation results further confirm that RAISE significantly outperforms other baselines in selecting logically relevant documents.
-5 Related Works
-LLMs for Scientific Reasoning.
-Recent works have shown that LLMs can be applied for challenging scientific reasoning tasks.
-Unlike other domains, scientific reasoning requires not only step-by-step thinking, but also knowledge of specialized terminology and adaptation to continually evolving knowledge.
-Due to this challenging nature, many benchmarks have been proposed recently to tackle scientific reasoning with LLMs Rein et al.
-(2024); Zhong et al.
-(2025); Zhang et al.
-(2025b); Wang et al.
-(2024a); Jiang et al.
-(2024).
-Many works enhance scientific reasoning capabilities of LLMs through domain-specific training Taylor et al.
-(2022); Prabhakar et al.
-(2025); Zhang et al.
-(2024b), step-by-step reasoning Rueda et al.
-(2025); Rein et al.
-(2024), or retrieval of external knowledge or tools Ma et al.
-(2024); Zhong et al.
-(2025); Wellawatte et al.
-(2025); Li et al.
-(2025).
-Unlike previous works, we focus on applying step-by-step document retrieval from in-the-wild corpus without assuming access to well-curated and domain-specific corpus.
-Step-wise Reasoning.
-A growing body of research has shown that decomposing complex problems into structured intermediate steps can enhance the reasoning abilities of LLMs.
-An influential early approach, Chain-of-Thought prompting Wei et al.
-(2022), introduced explicit, sequential reasoning steps, making the model’s thought process more transparent and coherent.
-This inspired methods such as Plan-and-Solve Wang et al.
-(2023), which emphasizes high-level planning before answering, and Step-Back Prompting Zheng et al.
-(2024), which encourages abstraction by prompting the model to reflect before solving.
-Least-to-Most prompting Zhou et al.
-(2023) extends this by breaking down tasks into simpler subproblems, solved in increasing order of difficulty.
-While prior work has focused on prompting strategies that help LLMs better use their internal reasoning capabilities, our work addresses a complementary challenge: enabling LLMs to retrieve and apply information from in-the-wild sources like Wikipedia, particularly during step-wise problem solving.
-We investigate how external evidence can be integrated at each step to improve reasoning beyond what internal knowledge alone can achieve.
-Retrieval Augmented Generation.
-Retrieval-Augmented Generation (RAG) Lewis et al.
-(2020b) was initially proposed to improve LLMs’ factual accuracy and knowledge by retrieving relevant external documents during generation Lewis et al.
-(2020b); Karpukhin et al.
-(2020); Izacard and Grave (2021); Khattab and Zaharia (2020).
-Recently, RAG has been extended for multi-hop reasoning, performing retrieval iteratively at multiple reasoning steps Liu et al.
-(2025); Zhao et al.
-(2024b); Zheng et al.
-(2024).
-In parallel, query reformulation and expansion techniques have been developed to enhance retrieval.
-Instead of using the original question, models generate enriched queries through prompting, such as intermediate answers or summaries.
-For example, HyDE Gao et al.
-(2023a) and CSQE Lei et al.
-(2024) demonstrate that carefully crafted queries greatly improve retrieval in complex, multi-step tasks.
-Building on this line of work, we redesign query expansion techniques with the specific goal of retrieving documents that contain the key logic or underlying principles required at each step of a step-wise reasoning process.
-This enables the model to supplement its limited internal knowledge with external sources, leading to more complete problem solving, especially in complex, multi-step tasks.
-6 Conclusion and Limitations
-We introduce RAISE, a step-by-step retrieval framework for scientific reasoning.
-We first decompose the problem into multiple subquestions and search queries, and then generate logical queries and retrieve logically relevant documents from in-the-wild corpus.
-We demonstrate the effectiveness of RAISE on three scientific reasoning benchmarks by comparing with various baselines.
-Our analysis shows that RAISE retrieves documents that are not only relevant in terms of the domain (e.g.
-definition of specificalized terminology) but also logically relevant documents for each subquestion, assisting the step-by-step reasoning process required for scientific reasoning.
-However, RAISE has some limitations and areas for future improvement.
-First, we used only DPR for retrieval.
-Incorporating recent models Gao et al.
-(2023b); Basu et al.
-(2024); Shi et al.
-(2023) may further improve performance.
-In addition, our study focuses on scientific reasoning.
-Applying RAISE to other domains such as legal, commonsense, or medical reasoning could help evaluate its generalizability.
-Moreover, RAISE requires multiple LLM calls during logical query generation.
-Reducing this cost through better prompting, caching, or lightweight planning could improve scalability and real-world usability.
-Appendix A RAISE Algorithm
-Input: Original question $x$, prompts $\mathcal{P}=\{p_{1},p_{2},p_{3},p_{4}\}$, corpus $\mathcal{C}$
-Output: Final answer $y$
-Step 1: Problem Decomposition
-Generate subquestions and initial queries:
-$\{(r_{i},q_{i})\}_{i=1}^{n}\sim P_{\theta}(\cdot\mid x,p_{1})$
-for $i=1$ to $n$ do
- Step 2: Logical Query Generation
-   Reformulate initial query:
-    $q_{i}^{*}\sim P_{\theta}(\cdot\mid r_{i},q_{i},p_{2})$
- Step 3: Knowledge Retrieval
-[…]
-    $D_{i}=\mathcal{R}(q_{i}^{*},\mathcal{C},k)$
- Step 4: Subquestion Answering
- if $i=1$ then
-       $a_{i}\sim P_{\theta}(\cdot\mid x,r_{1},D_{1},p_{3})$
- else
-       $a_{i}\sim P_{\theta}(\cdot\mid x,\{(r_{j},a_{j})\}_{j=1}^{i-1},r_{i},D_{i},p_{3})$
-   end if
-end for
-Step 5: Final Answer Composition
-Generate final answer using all subanswers:
-$y\sim P_{\theta}(\cdot\mid x,\{(r_{i},a_{i})\}_{i=1}^{n},p_{4})$
-Algorithm 1 RAISE Inference Procedure
-Appendix B Dataset Details
-GPQA
-Rein et al.
-(2024) This dataset consists of physics, biology, and chemistry questions written by domain experts.
-We use GPQA diamond subset, which consist of 198 high-quality questions selected based on human performance.
-Specifically, this subset includes questions that both experts answer correctly while the majority of non-experts fail to solve.
-Each question typically demands multi-step reasoning, precise formula manipulation, and access to external scientific facts (e.g., physical constants, definitions).
-Due to its alignment with our target setting, GPQA serves as the primary evaluation benchmark throughout our experiments.
-For GPQA, the original dataset does not include standardized multiple-choice labeled as (A), (B), (C) and (D).
-To ensure consistency during evaluation, we proprocessed each question by randomly shuffling the correct answer along with the three distractors, and assigning them uniformly to choice labels (A) through (D).
-SuperGPQA
-Team et al.
-(2025) SuperGPQA is a large-scale benchmark designed to evaluate graduate-level reasoning across 13 disciplines, 72 fields, and 285 graduate-level disciplines.
-In alignment with the scientific reasoning focus of our work, we select science and engineering domains for evaluation.
-Each domain is further divided by three difficulty levels(easy, medium, and hard).
-To reduce computational overhead while maintaining consistency, we randomly sample 199 questions per subset using a fixed seed (42).
-Specifically, our experiments include 199 examples each from science-hard, science-middle, and engineering-hard subsets.
-MMLU
-Hendrycks et al.
-(2021); Wang et al.
-(2024b) The MMLU benchmark covers a wide range of subjects across multiple domains.
-For out experiments, we focus on STEM and Professional categories.
-The STEM contains university-level science and engineering subjects such as college mathematics and computer science, while the Professional category covers specialized fields that typically require professional training or advanced education, including law, medicine, and chemistry.
-We specifically select three subsets: college chemistry from MMLU-STEM and professional chemistry and biology from MMLU-Pro.
-These subsets are chosen to evaluate our method’s ability to perform scientific reasoning in both academic and professional contexts involving complex domain knowledge.
-Appendix C Experiment Details
-C.1 Baseline Details
-CoT Wei et al.
-(2022); Kojima et al.
-(2023) We apply Chain-of-Thought prompting for direct reasoning, where the model is encouraged to explicitly generate intermediate reasoning steps through prompting (Think step by step).
-CoT+RAG
-Lewis et al.
-(2020b) We implement CoT+RAG by combining Chain-of-Thought prompting with retrieval, where the model is prompted to solve the problem step-by-step while also leveraging external knowledge.
-Specifically, we provide the model with a CoT-style prompt encouraging step-by-step reasoning, alongside the original question and documents retrieved using the original question as the search query.
-Least-to-Most
-Zhou et al.
-(2023) Least-to-Most is a decomposed reasoning strategy that breaks down a complex problem into a sequence of simpler subquestions, which are then solved sequentially without retrieval augmentation.
-This subquestion decomposition pipeline serves as the foundational structure for other decomposed reasoning methods as well.
-Step-Back
-(Decomposed reasoning) Zheng et al.
-(2024) We implement Step-Back for decomposed reasoning by applying the Step-Back prompting method to each subquestion in a decomposed reasoning framework.
-While the original Step-Back paper does not cover the application of this method to decomposed subquestions, we extend it for a fair comparison with our approach.
-Specifically, after decomposing the original question into subquestions, we use the Step-Back prompting strategy to extract a high-level principle for each subquestion, and then provide the subquestion along with its corresponding principle to guide the model’s reasoning.
-Least-to-Most+RAG
-(Decomposed reasoning with RAG) Liu et al.
-(2025) We implement RAG by first decomposing the original problem into subquestions and then retrieving documents using each subquestion as a query.
-The retrieved documents are provided to the model along with the corresponding subquestion to support its reasoning.
-Step-Back+RAG
-(Decomposed reasoning with RAG) Zheng et al.
-(2024) We extend the Step-Back prompting strategy to a retrieval-augmented setting for fair comparison with our method.
-After decomposing the original question into subquestions, we generate a principle abstraction for each subquestion using Step-Back prompting, and use it as a query to retrieve evidence.
-The retrieved documents are then provided alongside the original subquestion to guide the model’s reasoning.
-HyDE
-(Decomposed reasoning with RAG) Gao et al.
-(2023a) We apply the HyDE approach to each subquestion in a decomposed reasoning framework.
-For each subquestion, the model first generates a hypothetical answer, which is then used as a query to retrieve supporting documents.
-The retrieved evidence, together with the subquestion, is provided to the model to support step-by-step reasoning.
-C.2 Retriever Configuration
-We use the pre-trained DPR encoder from the ’facebook/dpr-question_encoder-single-nq-base’ model Karpukhin et al.
-(2020), which is a BERT-based encoder trained for open-domain question answering.
-This encoder is trained on the Natural Question (NQ) dataset Kwiatkowski et al.
-(2019) and is designed to map questions into 768-dimensional dense vector representations for retrieval.
-For the retrieval corpus, we use the preprocessed Wikipedia passages provided by ’facebook/wiki_dpr’ Karpukhin et al.
-(2020), a corpus widely used to evaluate DPR-based retrieval models.
-This corpus is constructed from the December 20, 2018 Wikipedia dump, where each article is split into multiple, disjoint text blocks of 100 words, resulting in approximately 21 million passages.
-Each passage is accompanied by the title of the wikipedia page it comes from along with DPR embedding.
-To enable efficient retrieval over the passage embeddings, we use an exact FAISS index.
-FAISS (Facebook AI Similarity Search) Johnson et al.
-(2019); Douze et al.
-(2024) is widely used library for fast similarity search over dense vectors.
-Throughout all experiment, we retrieve top-10 documents per query.
-To reduce the impact of potentially irrelevant documents by DPR, we apply a similarity threshold $T$ in RAISE.
-Specifically, we discard any retrieved passage whose DPR similarity score falls below $T$.
-DPR similarity is computed as the inner product between L2-normalized query and passage embeddings.
-Higher scores indicate greater semantic similarity, with values closer to 1 representing stronger alignment between the query and passage.
-We set $T=0.84$ for GPQA, SuperGPQA, and MMLU-Pro, which are composed of more challenging reasoning problems.
-For MMLU-STEM (college chemistry), we use a slightly lower threshold of $T=0.80$, considering that the questions are generally simpler than those in other datasets.
-Appendix D Additional Results
-D.1 Further Analysis on GPQA
-Figure 5: Domain-wise accuracy comparison between CoT and RAISE on the GPQA Diamond subset
+Since our domain includes complex, expert-level questions, and LLM-based evaluations may overlook domain-specific reasoning and often rely on surface-level features, we supplemented our analysis with a small-scale human evaluation of 20 subquestion–document pairs.
+Each pair was assessed by at least three annotators, including Ph.D.
+students and a faculty member in chemistry, with the method provenance concealed to maintain objectivity.
+As also discussed in the LLM-as-a-judge results, the human evaluation indicates that RAISE produces significantly fewer irrelevant documents compared to all other methods, while achieving the highest proportion of logically relevant documents.
+Although limited in scale due to time and cost constraints, we believe this evaluation provides meaningful human validation of RAISE’s effectiveness and serves as a valuable complement to the LLM-based assessments.
+C.4 Further Analysis on GPQA
+Figure 7: Domain-wise accuracy comparison between CoT and RAISE on the GPQA Diamond subset.
 Figure 7 shows the domain-wise accuracy on the GPQA Diamond dataset.
@@
 These results demonstrate RAISE’s robustness and its ability to generalize across diverse areas of graduate-level scientific reasoning.
-D.2 Human Evaluation of Retrieved Documents
-Figure 6: Human Evaluation of the Logical Relevance of Retrieved Documents Aligned with the results from the LLM-as-a-judge evaluation of logical relevancy, RAISE shows a higher proportion of logically relevant documents and a lower proportion of irrelevant or superficially relevant ones.
-Previously, in Section 4.3.2, we evaluated the logical relevancy of documents retrieved by the baseline methods and RAISE using the LLM-as-a-judge approach.
-However, since our domain includes complex, expert-level questions, and LLM-based evaluations may overlook domain-specific reasoning and often rely on surface-level features, we supplemented our analysis with a small-scale human evaluation of 20 subquestion–document pairs.
-Each pair was assessed by at least three annotators, including Ph.D.
-students and a faculty member in chemistry, with the method provenance concealed to maintain objectivity.
-The results are presented in Figure 6.
-As also discussed in the LLM-as-a-judge results, the human evaluation indicates that RAISE produces significantly fewer irrelevant documents compared to all other methods, while achieving the highest proportion of logically relevant documents.
-Although limited in scale due to time and cost constraints, we believe this evaluation provides meaningful human validation of RAISE’s effectiveness and serves as a valuable complement to the LLM-based assessments.
+Appendix D Related Works
+LLMs for Scientific Reasoning.
+Recent works have shown that LLMs can be applied for challenging scientific reasoning tasks.
+Unlike other domains, scientific reasoning requires not only step-by-step thinking, but also knowledge of specialized terminology and adaptation to continually evolving knowledge.
+Due to this challenging nature, many benchmarks have been proposed recently to tackle scientific reasoning with LLMs [29, 50, 46, 35, 10].
+Many works enhance scientific reasoning capabilities of LLMs through domain-specific training [31, 27, 43], step-by-step reasoning [30, 29], or retrieval of external knowledge or tools [24, 50, 38, 22].
+Unlike previous works, we focus on applying step-by-step document retrieval from in-the-wild corpus without assuming access to well-curated and domain-specific corpus.
+Step-wise Reasoning.
+A growing body of research has shown that decomposing complex problems into structured intermediate steps can enhance the reasoning abilities of LLMs.
+An influential early approach, Chain-of-Thought prompting [37], introduced explicit, sequential reasoning steps, making the model’s thought process more transparent and coherent.
+This inspired methods such as Plan-and-Solve [33], which emphasizes high-level planning before answering, and Step-Back Prompting [49], which encourages abstraction by prompting the model to reflect before solving.
+Least-to-Most prompting [51] extends this by breaking down tasks into simpler subproblems, solved in increasing order of difficulty.
+While prior work has focused on prompting strategies that help LLMs better use their internal reasoning capabilities, our work addresses a complementary challenge: enabling LLMs to retrieve and apply information from in-the-wild sources like Wikipedia, particularly during step-wise problem solving.
+We investigate how external evidence can be integrated at each step to improve reasoning beyond what internal knowledge alone can achieve.
+Retrieval Augmented Generation.
+Retrieval-Augmented Generation (RAG) [21] was initially proposed to improve LLMs’ factual accuracy and knowledge by retrieving relevant external documents during generation [21, 14, 8, 15].
+Recently, RAG has been extended for multi-hop reasoning, performing retrieval iteratively at multiple reasoning steps [23, 48, 49].
+In parallel, query reformulation and expansion techniques have been developed to enhance retrieval.
+Instead of using the original question, models generate enriched queries through prompting, such as intermediate answers or summaries.
+For example, HyDE [3] and CSQE [19] demonstrate that carefully crafted queries greatly improve retrieval in complex, multi-step tasks.
+Building on this line of work, we redesign query expansion techniques with the specific goal of retrieving documents that contain the key logic or underlying principles required at each step of a step-wise reasoning process.
+This enables the model to supplement its limited internal knowledge with external sources, leading to more complete problem solving, especially in complex, multi-step tasks.
 Appendix E Prompts